### v1

# Weighted wave envelope estimates for the parabola

Abstract.
In this paper, we extend Fefferman’s classical square function estimate for the parabola to a weighted setting. Our weighted square function estimate is derived from a weighted wave envelope estimate for the parabola. The bounds are formulated in terms of families of multiscale tubes together with weight parameters that quantify the distribution of the weight. As an application, we obtain some weighted $L^{p}$-estimates for a class of Fourier multiplier operators and for solutions to free Schrödinger equation.
Key words and phrases:
square function estimates, wave envelope estimates, weighted estimates
2020 Mathematics Subject Classification:
42B15, 42B25
1. Introduction
The paper is concerned with weighted square function estimates for the parabola and some of its applications. Let $\mathcal{P}$ denote the truncated parabola
$\mathcal{P}=\{(t,t^{2})\in\mathbb{R}^{2}:|t|\leq 1\}$
and $N_{R^{-1}}\mathcal{P}=\{(t,t^{2}+\eta)\in\mathbb{R}^{2}:|t|\leq 1,\;|\eta|\leq R^{-1}\}$ denote its $R^{-1}$-neighborhood for a large $R\geq 1$. We consider the canonical covering of $N_{R^{-1}}\mathcal{P}$ by finitely overlapping parallelograms $\theta$ of dimensions $R^{-1/2}\times R^{-1}$. Given a function $f$ whose Fourier transform is supported on $N_{R^{-1}}\mathcal{P}$, we decompose $f=\sum_{\theta}f_{\theta}$, where $\widehat{f}_{\theta}$ is supported on $\theta$. This can be done, for example, by using a smooth partition of unity subordinate to a covering of the interval [1,1][-1,1] by finitely overlapping intervals of length $\sim R^{-1/2}$ (see e.g. the proof of Theorem 2.1).
By Plancherel’s theorem, these functions $\{f_{\theta}\}$ are orthogonal on $L^{2}(\mathbb{R}^{2})$: $\|f\|_{L^{2}}^{2}\leq C\sum_{\theta}\|f_{\theta}\|_{L^{2}}^{2}$. Moreover, the family exhibits certain $L^{p}$ orthogonality due to the curvature properties of the parabola for some $p$ larger than $2$. For instance, the classical square function estimate for the parabola (see [9]) states that
(1) $\|f\|_{L^{4}(\mathbb{R}^{2})}\leq C\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{4}(\mathbb{R}^{2})}.$
This inequality relies on the geometric observation by Fefferman [14] that the algebraic differences $\theta-\theta^{\prime}$ overlap only finitely often as $\theta\neq\theta^{\prime}$ vary. See also [17, 30, 20] for extensions to non-degenerate curves in higher dimensions. Square function estimates of the form (1) have several important applications in harmonic analysis. The sharp square function estimate (1) is known to imply sharp results for the Kakeya maximal function, the Bochner-Riesz multipliers, the Fourier restriction operator, and local smoothing estimates for the Schrödinger equation; see [6, 41] and references therein. In higher dimensions, it is conjectured that (1) holds with $L^{4}(\mathbb{R}^{2})$ replaced by $L^{\frac{2d}{d-1}}(\mathbb{R}^{d})$, which remains wide open.
Let $H:\mathbb{R}^{2}\to[0,\infty)$ be a bounded function on $\mathbb{R}^{2}$. The main goal of this paper is to establish weighted square function estimates of the form
$\|f\|_{L^{p}(Hdx)}\leq C_{p,H}(R)\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}$
for $2\leq p\leq 4$ and to explore some of its consequences. To describe the constant $C_{p,H}(R)$, we need to introduce a family of tubes originating from a multiscale analysis. Let $s$ be a dyadic number in the range $R^{-1/2}\leq s\leq 1$. At each scale $s$, we cover $N_{s^{2}}\mathcal{P}$ by canonical blocks $\{\tau\}$ of dimension $s\times s^{2}$ and use $|\tau|=s$ to denote the scale. For the smallest scale $s=R^{-1/2}$, these blocks are just $\{\theta\}$.
We fix a dyadic $s\in[R^{-1/2},1]$. For each $\tau$ with $|\tau|=s$, we consider a linear transform $L_{\tau}$ determined by the parabolic rescaling (see (38)) for which $L_{\tau}([-\frac{1}{2},\frac{1}{2}]^{2})$ is a parallelepiped dual to $\tau$ of dimensions $s^{-1}\times s^{-2}$ and orthogonal to $\tau$. Let $\mathbb{T}_{\tau}$ denote the tiling of $\mathbb{R}^{2}$ by translates of the dual parallelepiped:
(2) $\mathbb{T}_{\tau}=\{L_{\tau}(z+q):z\in\mathbb{Z}^{2}\},\;\;q=[-1/2,1/2]^{2}.$
Next, we consider the tiling of $\mathbb{R}^{2}$ by the dilated family of tubes
$\mathbb{U}_{\tau}=\{Rs^{2}\cdot T:T\in\mathbb{T}_{\tau}\}.$
Each $U\in\mathbb{U}_{\tau}$ is thus a parallelepiped of dimensions $Rs\times R$. Let $\mathbb{U}$ denote the union of $\mathbb{U}_{\tau}$ for all $\tau$ ranging over all dyadic scales $R^{-1/2}\leq s\leq 1$. For a given $U\in\mathbb{U}$, we let $\tau(U)$ denote the $\tau$ such that $U\in\mathbb{U}_{\tau}$.
[…]
Figure 1. $U\in\mathbb{U}_{\tau}$, $T\in\mathbb{T}_{\tau}$ such that $T\subset U$ for $|\tau|=s$.
Given $U\in\mathbb{U}$, we define
(3) $\kappa_{p,H}(U)=\max_{\begin{subarray}{c}T\in\mathbb{T}_{\tau(U)}:\\ T\subset U\end{subarray}}\left(\frac{H(T)}{|T|}\right)^{\frac{1}{4}}\left(\frac{H(U)}{|U|}\right)^{\frac{1}{p}-\frac{1}{4}},$
where we write $H(E):=\int_{E}H$ for a measurable set $E\subset\mathbb{R}^{2}$. We are now ready to state our weighted square function estimates.
Theorem 1.1.
Let $2\leq p\leq 4$ and $H:\mathbb{R}^{2}\to[0,1]$ be a function. For any function $f$ whose Fourier transform is supported on $N_{R^{-1}}(\mathcal{P})$, we have,
(4) $\|f\|_{L^{p}(Hdx)}\lessapprox\big(\max_{U\in\mathbb{U}}\kappa_{p,H}(U)+R^{-100}\big)\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}.$
Here we mean by $A\lessapprox B$ an inequality of the form $A\leq C_{p}(\log R)^{O(1)}B$. The term $R^{-100}$ on the right-hand side of (4) is negligible for interesting weights $H$. For instance, we have $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)\gg R^{-100}$ whenever $H$ is the characteristic function of a union of unit balls. We also use the notation $A\lesssim B$ to denote $A\leq CB$ with an absolute constant $C>0$, possibly depending on parameters such as $p$ and $\alpha$.
Remark 1.2.
Theorem 1.1 (and Theorem 1.4 to be stated) remains valid when the parabola $\mathcal{P}$ is replaced by a small perturbation of $\mathcal{P}$ for which the bilinear restriction estimate (see Theorem 4.4) is valid. In particular, it holds for any function whose Fourier transform is supported in a small neighborhood of the unit circle under a corresponding modification in the collections $\mathbb{T}_{\tau}$ and $\mathbb{U}_{\tau}$ outlined before the statement of Theorem 2.1 below. In addition, (4) holds for all non-negative $H\in L^{\infty}(\mathbb{R}^{2})$ with $R^{-100}$ replaced by $R^{-100}\|H\|_{L^{\infty}}^{1/p}$ by homogeneity and the fact that $\kappa_{p,cH}(U)=c^{1/p}\kappa_{p,H}(U)$ for any $U\in\mathbb{U}$ and constant $c>0$.
For $H\equiv 1$, we have $\kappa_{p,H}(U)=1$ for any $U\in\mathbb{U}$. Thus, when $p=4$, Theorem 1.1 essentially recovers the classical square function estimate (1). For $2\leq p<4$, the $H\equiv 1$ case of Theorem 1.1 recovers square function estimates due to Gan [15], where more general small cap square function estimates are established. Our weighted square function estimates are inspired by weighted decoupling inequalities for the paraboloids, which have been extensively studied in recent years and applied to problems such as the Falconer distance set conjecture and Bochner-Riesz means; see, e.g., [18, 11, 16, 26] and references therein.
We compute the constant $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)$ for $\alpha$-dimensional weights.
Example 1 ($\alpha$-dimensional weights).
Let $0\leq\alpha\leq 2$. Suppose that $H:\mathbb{R}^{2}\to[0,1]$ is $\alpha$-dimensional in the sense that
$\langle H\rangle_{\alpha}:=\sup_{(z,\rho)\in\mathbb{R}^{2}\times[1,\infty]}\rho^{-\alpha}H(B_{\rho}(z))\lesssim 1.$
Here $B_{\rho}(z)$ denotes the ball of radius $\rho$ centered at $z$ (and we simply write $B_{\rho}$ when centered at the origin). Then
(5) $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)\;\lesssim\;R^{-(2-\alpha)\left(\frac{1}{p}-\frac{1}{4}\right)}.$
To see this, let $T\in\mathbb{T}_{\tau}$ and $U\in\mathbb{U}_{\tau}$ for some $|\tau|=s$. Since $T$ and $U$ are covered by $O(s^{-1})$ balls of radius $s^{-1}$ and $Rs$, respectively, we have
$H(T)$ $\lesssim s^{-1}s^{-\alpha},$
$H(U)$ $\lesssim s^{-1}(Rs)^{\alpha}.$
On the other hand, $|T|\sim s^{-3}$ and $|U|\sim R^{2}s$. Therefore,
$\max_{U\in\mathbb{U}}\kappa_{p,H}(U)\;\lesssim\;\max_{R^{-1/2}\leq s\leq 1}(s^{2-\alpha})^{\frac{1}{4}}\big((Rs)^{-(2-\alpha)}\big)^{\frac{1}{p}-\frac{1}{4}},$
and the maximum is attained at the scale $s=1$, which yields (5).
Example 2 (Unit ball).
We examine the sharpness of Theorem 1.1 for the weight $H=1_{B_{1}}$. Since $H$ is $\alpha$-dimensional for every $\alpha\in[0,2]$, Theorem 1.1 together with (5) shows that
$\|f\|_{L^{p}(B_{1})}\lessapprox R^{-2(\frac{1}{p}-\frac{1}{4})}\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}.$
This estimate is essentially sharp for all $2\leq p\leq 4$. Indeed, let $\widehat{f_{\theta}}$ be an $L^{1}$-normalized smooth bump function supported on $\theta$. In this case, $|\sum_{\theta}f_{\theta}(x)|\gtrsim\#\{\theta\}$ for $x\in B_{c}$ for a sufficiently small $c>0$ and $|f_{\theta}|$ decays rapidly away from the tube $\theta^{*}$ dual to $\theta$ centered at the origin, implying
$\|f\|_{L^{p}(B_{1})}\gtrsim R^{\frac{1}{2}}\;\text{ and }\;\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}\sim R^{\frac{2}{p}}.$
Remark 1.3.
Fix $\alpha\in[1,2]$. Let $\sigma$ denote the infimum of exponents for which the bound
$\|f\|_{L^{p}(Hdx)}\lessapprox R^{\sigma}\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{\frac{1}{2}}\big\|_{L^{p}(\mathbb{R}^{2})}$
holds for any weight $H$ with $\langle H\rangle_{\alpha}\leq 1$. Theorem 1.1 and (5) yield the upper bound
(6) $\sigma\leq-(2-\alpha)\left(\frac{1}{p}-\frac{1}{4}\right).$
When $p=2$ or $p=4$, the upper bound matches with the lower bound
(7) $\sigma\geq\max\Big(-2\big(\frac{1}{p}-\frac{1}{4}\big),~-\frac{2-\alpha}{2p},~-(2-\alpha)\big(\frac{1}{p}-\frac{1}{6}\big)\Big).$
The first lower bound follows from 2. For the second lower bound, fix $\theta$ and take $f=f_{\theta}$ as in 2, and set $H=R^{\frac{\alpha-2}{2}}1_{\theta^{*}}$. A direct computation shows $\langle H\rangle_{\alpha}\lesssim 1$, which yields the second lower bound. The third lower bound can be obtained by using a special solution to Schrödinger equation studied by Barceló, Bennett, Carbery, Ruiz, and Vilela [1]; see Section 3.1.
On the other hand, there are gaps between (6) and (7) for intermediate $2<p<4$. Nevertheless, we will show that, for each $\alpha\in(1,2)$, there exists an $\alpha$-dimensional weight for which Theorem 1.1 gives sharp $L^{p}$ weighted square function estimates when $2\leq p\leq 4/(3-\alpha)$ or $p=4$; see Section 3.2.
We present three consequences of Theorem 1.1 in Section 2.
$(i)$ Weighted $L^{p}$ bounds for Fourier multipliers supported on a small neighborhood of the unit circle.
$(ii)$ Weighted and frequency–localized $L^{p}$ bounds for the one–dimensional Schrödinger propagator.
$(iii)$ Local smoothing estimates for the Schrödinger equation with respect to fractal measures satisfying parabolic or Euclidean ball conditions.
In cases $(i)$ and $(ii)$, the dependence on the weight is quantified by $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)$, which extends classical unweighted estimates.
Theorem 1.1 is a consequence of a weighted $L^{p}$ wave envelope estimate for the parabola.
Theorem 1.4.
Let $2\leq p\leq 4$ and $H:\mathbb{R}^{2}\to\{0\}\cup[R^{-400},1]$ be a weight. If $\widehat{f}$ is supported on $N_{R^{-1}}(\mathcal{P})$, we have
(8) $\|f\|_{L^{p}(Hdx)}^{p}\lessapprox\sum_{R^{-1/2}\leq s\leq 1}\sum_{|\tau|=s}\sum_{U\in\mathbb{U}_{\tau}}\kappa_{p,H}(U)^{p}|U|^{1-\frac{p}{2}}\big\|(\sum_{\theta\subset\tau}|f_{\theta}|^{2})^{1/2}\big\|_{L^{2}(w_{U})}^{p}.$
Here $w_{U}$ denotes an $L^{\infty}$-normalized weight which decays rapidly away from $U$.
The estimate (8) is sharp for the unit ball example in 2, where the term $s=1$ on the right-hand side of (8) dominates. We present an example where $\|f\|_{L^{p}(Hdx)}^{p}$ essentially matches the contribution from $s=R^{-1/2}$.
Example 3.
Let $Y\subset\mathbb{R}^{2}$ and $H=1_{Y}$. Consider $f=\sum_{\theta}f_{\theta}$ such that $\{f_{\theta}\}$ have essentially disjoint supports on $Y$. Then
$\|f\|_{L^{p}(Y)}^{p}\sim\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}=\sum_{\theta}\sum_{U\in\mathbb{U}_{\theta}}\|f_{\theta}\|_{L^{p}(U\cap Y)}^{p}.$
We further assume that $|f_{\theta}|$ is essentially constant on each $U\in\mathbb{U}_{\theta}$, which is natural in view of the uncertainty principle. Then
(9) $\|f_{\theta}\|_{L^{p}(U\cap Y)}^{p}\sim\frac{|U\cap Y|}{|U|}\|f_{\theta}\|_{L^{p}(U)}^{p}\sim\frac{|U\cap Y|}{|U|}|U|^{1-\frac{p}{2}}\|f_{\theta}\|_{L^{2}(U)}^{p}.$
Hence,
$\|f\|_{L^{p}(Y)}^{p}\;\sim\;\sum_{|\theta|=R^{-1/2}}\sum_{U\in\mathbb{U}_{\theta}}\frac{|U\cap Y|}{|U|}\,|U|^{1-\frac{p}{2}}\,\big\|f_{\theta}\big\|_{L^{2}(U)}^{p}.$
We note that $\mathbb{U}_{\theta}=\mathbb{T}_{\theta}$ forms an identical tiling of $\mathbb{R}^{2}$ by parallelepipeds of dimensions $R^{1/2}\times R$. Therefore,
(10) $\kappa_{p,H}(U)^{p}=\frac{|U\cap Y|}{|U|},\;\;U\in\mathbb{U}_{\theta}.$
Thus, $\|f\|_{L^{p}(Y)}^{p}$ is comparable to the $s=R^{-1/2}$ term on the right-hand side of (8).
Wave envelope estimates, namely estimates of the form (8) with $H\equiv 1$, were first developed in the breakthrough work of Guth, Wang, and Zhang [21] for the cone
$\Gamma=\{\xi_{1}^{2}+\xi_{2}^{2}=\xi_{3}^{2},~1/2\leq\xi_{3}\leq 2\}\subset\mathbb{R}^{3}.$
Their wave envelope estimate for $\Gamma$ implies, among other consequences, the sharp $L^{4}$ square function estimate for $\Gamma$ and the sharp local smoothing estimate for the wave equation in 2+1 dimension. The $p=4$ and $H\equiv 1$ case of Theorem 1.4 recovers the $L^{4}$ wave envelope estimate [19, Equation (8)], which is implicit in [21]. We note that [19] established more refined versions of the $L^{4}$ wave envelope estimates, termed amplitude-dependent wave envelope estimates, for both the parabola $\mathcal{P}$ and the cone $\Gamma$.
For the proof of Theorem 1.4, the classical approach used to establish the square function estimate (1) is not applicable, as it relies critically on the even exponent $4$ and Plancherel’s theorem, neither of which extend to weighted settings or general exponents. Instead, we adopt a more robust strategy used by Gan [15] for proving small cap square function estimates. This method employs a multiscale bilinear reduction argument from [5, 10], together with the bilinear restriction theorem (see e.g. [39]). One of main contributions of the present paper is an extension of the method to the weighted setting that effectively exploits the presence of the weights without imposing any additional assumptions.
Organization of the paper
In Section 2, we deduce Theorem 1.1 from Theorem 1.4 and discuss applications of Theorem 1.1. In Section 3, we present additional examples related to the sharpness of Theorem 1.1 for $\alpha$-dimensional weights. In Section 4, we establish Theorem 1.4 using the bilinear restriction theorem. In Section 5, we prove fractal local smoothing estimates to be stated in Section 2. In Section 6, we look at examples and derive necessary conditions for 2.3, Theorem 2.4 and Theorem 2.5. Finally, in Appendix A, we give a proof of Lemma 4.1, a multiscale broad-narrow decomposition.
Notations
We summarize here the notations that will be used frequently throughout the paper.
$\bullet$ We write $A\lessapprox B$ to denote an inequality of the form $A\leq C_{p}(\log R)^{O(1)}B$ for $R>1$ where $C_{p}$ is a constant depending only on $p$.
$\bullet$ We denote by $B_{\rho}(z)$ the ball of radius $\rho$ centered at $z\in\mathbb{R}^{2}$ and we simply write $B_{\rho}$ when centered at the origin.
$\bullet$ For a Borel measure $\mu$ on $\mathbb{R}^{2}$, we set ${\langle\mu\rangle_{\alpha}:=\sup_{(z,\rho)\in\mathbb{R}^{2}\times[1,\infty)}\rho^{-\alpha}\mu(B_{\rho}(z))}$ and define $\mu]_{\alpha}$ similarly except that the supremum is taken over $\rho>0$. Analogous conventions apply for other related quantities.
2. Proof of Theorem 1.1 and some applications
We begin by deriving Theorem 1.1 from the weighted envelope estimate in Theorem 1.4. Then we turn to some applications of Theorem 1.1.
2.1. Weighted wave envelope estimates imply weighted square function estimates
In this section, we prove that Theorem 1.4 implies Theorem 1.1.
Let $2\leq p\leq 4$. We first verify (4) for weights $H:\mathbb{R}^{2}\to[R^{-400},1]$. By Hölder’s inequality, Theorem 1.4 yields
$\|f\|_{L^{p}(Hdx)}^{p}\lessapprox\sum_{R^{-1/2}\leq s\leq 1}\sum_{|\tau|=s}\sum_{U\in\mathbb{U}_{\tau}}\kappa_{p,H}(U)^{p}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(w_{U})}^{p}.$
After dominating $\kappa_{p,H}(U)^{p}$ by $\sup_{U\in\mathbb{U}}\kappa_{p,H}(U)^{p}$, we sum over all $U$. This yields, for each $s$,
$\sum_{|\tau|=s}\sum_{U\in\mathbb{U}_{\tau}}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(w_{U})}^{p}\lesssim\sum_{|\tau|=s}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}}^{p}\leq\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}}^{p}.$
For the last inequality, we use embedding $\ell^{2}\subset\ell^{p}$ for $p\geq 2$. Since $s$ ranges over dyadic numbers in $R^{-\frac{1}{2}},1]$, this gives Theorem 1.1 when $H:\mathbb{R}^{2}\to[R^{-400},1]$.
For the case $H:\mathbb{R}^{2}\to[0,1]$, we decompose $H=H_{1}+H_{2}$, where $0\leq H_{1}\leq R^{-400}$ and $R^{-400}\leq H_{2}\leq 1$. For $H_{2}$, we have already obtained a bound which involves $\sup_{U\in\mathbb{U}}\kappa_{p,H}(U)$. For $H_{1}$, we use the unweighted case ($H\equiv 1$) to get
$\|f\|_{L^{p}(H_{1})}\leq R^{-100}\|f\|_{L^{p}(\mathbb{R}^{2})}\lessapprox R^{-100}\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}}.$
Combining these estimates yields Theorem 1.1. ∎
2.2. Weighted estimates for a radial Fourier multiplier
Let $\psi$ be a smooth bump function supported on [1,1][-1,1]. We consider the Fourier multiplier transformation $S_{R}$ defined by
$\widehat{S_{R}f}(\xi)=\psi(R(1-|\xi|))\widehat{f}(\xi),$
which plays a critical role in the theory of Bochner-Riesz means. It is well-known that
(11) $\|S_{R}f\|_{L^{p}(\mathbb{R}^{2})}\lessapprox\|f\|_{L^{p}(\mathbb{R}^{2})},\;\;2\leq p\leq 4,$
which follows from an interpolation of the trivial $L^{2}$-bound and the sharp $L^{4}$-bound due to Córdoba [9] which relies on the square function estimate (1) and bounds for the Nikodym maximal function.
We present a weighted version of (11). For each dyadic scale $R^{-1/2}\leq s\leq 1$, we cover $N_{R^{-1}}\mathbb{S}^{1}$ by finitely overlapping rectangles $\tau$ of dimensions $s\times s^{2}$ and define $\mathbb{T}_{\tau}$, $\mathbb{U}_{\tau}$ and $\mathbb{U}$, accordingly. With this minor modification in mind, we obtain the following.
Theorem 2.1.
Let $2\leq p\leq 4$ and $H:\mathbb{R}^{2}\to[0,1]$ be a function. Then
$\|S_{R}f\|_{L^{p}(Hdx)}\lessapprox\big(\max_{U\in\mathbb{U}}\kappa_{p,H}(U)+R^{-100}\big)\|f\|_{L^{p}(\mathbb{R}^{2})}.$
Proof.
The proof is essentially the same as the proof of (11) by Córdoba [9], so we only sketch the argument. We divide $\mathbb{R}^{2}$ into four sectors by lines $y=\pm x$. Without loss of generality, we may replace $S_{R}$ by a smooth frequency projection to the part of $N_{R^{-1}}\mathbb{S}^{1}$ contained in one of the four sectors which includes the point (0,1)(0,-1).
Next, we cover [1,1][-1,1] by finitely overlapping intervals $I$ of length $\sim R^{-1/2}$, and let $\{\chi_{I}\}$ be a smooth partition of unity adapted to this covering. Then we have $S_{R}f=\sum_{I}S_{R}f_{I}$, where $\widehat{f_{I}}(\xi_{1},\xi_{2})=\chi_{I}(\xi_{1})\widehat{f}(\xi_{1},\xi_{2})$. By Theorem 1.1, we have
$\|S_{R}f\|_{L^{p}(Hdx)}\lessapprox\big(\max_{U\in\mathbb{U}}\kappa_{p,H}(U)+R^{-100}\big)\big\|\big(\sum_{I}|S_{R}f_{I}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}.$
By duality and the boundedness of the Nikodym maximal function, for any $2\leq p\leq 4$,
$\big\|\big(\sum_{I}|S_{R}f_{I}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}\lessapprox\big\|\big(\sum_{I}|f_{I}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}.$
Finally, by the Littlewood-Paley inequality for equally spaced intervals, we have
(12) $\big\|\big(\sum_{I}|f_{I}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}\lesssim\|f\|_{L^{p}(\mathbb{R}^{2})},\;\;\;p\geq 2,$
which completes the proof. ∎
2.3. Weighted estimates for the Schrödinger equation
Let
$e^{it\partial_{x}^{2}}f(x)=(2\pi)^{-1}\int e^{ix\xi}e^{it\xi^{2}}\widehat{f}(\xi)\,d\xi$
denote the solution to the free Schrödinger equation
$\begin{cases}i\partial_{t}u=\partial_{x}^{2}u,&(x,t)\in\mathbb{R}\times\mathbb{R}\\ u(x,0)=f(x),&x\in\mathbb{R}.\end{cases}$
Let $\eta\in C^{\infty}_{c}(\mathbb{R})$ be such that $\widehat{\eta}$ is compactly supported on [1,1][-1,1] and $|\eta(t)|\sim 1$ on [1,1][-1,1]. We define the operator $\mathbf{U}_{R}$ by
(13) $\mathbf{U}_{R}f(x,t)=\eta(R^{-1}t)e^{it\partial_{x}^{2}}f(x).$
If $\widehat{f}$ is supported on [1,1][-1,1], then $\widehat{\mathbf{U}_{R}f}$ is supported on $N_{R^{-1}}\mathcal{P}$.
As another consequence of Theorem 1.1, we establish a weighted estimate for the Schrödinger propagator.
Theorem 2.2.
Let $2\leq p\leq 4$ and $H:\mathbb{R}^{2}\to[0,1]$ be a function. For any function $f$ whose Fourier transform is supported on [1,1][-1,1], we have
$\|\mathbf{U}_{R}f\|_{L^{p}(\mathbb{R}^{2},Hdxdt)}\lessapprox\big(\max_{U\in\mathbb{U}}\kappa_{p,H}(U)+R^{-100}\big)R^{\frac{1}{p}}\|f\|_{L^{p}(\mathbb{R})}.$
Proof.
Let $\{\chi_{I}\}$ be the smooth partition of unity given in Theorem 2.1. We decompose $f=\sum_{I}f_{I}$, where $\widehat{f_{I}}=\chi_{I}\widehat{f}$. By Theorem 1.1, we have
$\|\mathbf{U}_{R}f\|_{L^{p}(\mathbb{R}^{2},Hdxdt)}\lessapprox\big(\max_{U\in\mathbb{U}}\kappa_{p,H}(U)+R^{-100}\big)\big\|\big(\sum_{I}|\mathbf{U}_{R}f_{I}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}.$
Thus, it suffices to verify that
(14) $\Big\|\big(\sum_{I}|\mathbf{U}_{R}f_{I}|^{2}\big)^{1/2}\Big\|_{L^{p}(\mathbb{R}^{2})}\lessapprox R^{\frac{1}{p}}\|f\|_{L^{p}(\mathbb{R}^{2})},\;\;2\leq p\leq 4.$
A detailed proof of (14) can be found in a note by Yung [41, Proof of Theorem 2]. It follows from a standard duality argument similar to the one used in [9] and [33]. For completeness, we sketch the argument in Section 5.3. ∎
As a corollary, we state a special case for $\alpha$-dimensional measures. For $0<\alpha<1$ and a measure $\mu$ defined on $\mathbb{R}^{2}$, we set
(15) $\langle\mu\rangle_{\alpha}:=\sup_{(z,\rho)\in\mathbb{R}^{2}\times[1,\infty)}\rho^{-\alpha}\mu(B_{\rho}(z)).$
Corollary 2.3.
Let $2\leq p\leq 4$ and $0\leq\alpha\leq 2$. For any function $f$ whose Fourier transform is supported on [1,1][-1,1] and $\mu$ satisfying $\langle\mu\rangle_{\alpha}\leq 1$, we have
$\big\|\mathbf{U}_{R}f\big\|_{L^{p}(\mathbb{R}\times[0,R],\mu)}\lessapprox R^{\frac{1}{p}-(2-\alpha)\big(\frac{1}{p}-\frac{1}{4}\big)}\|f\|_{L^{p}(\mathbb{R})}.$
Proof.
Recall (13). By the Fourier localization property of $\mathbf{U}_{R}f$, we may write $\mathbf{U}_{R}f=\mathbf{U}_{R}f*\varphi$ for a Schwartz function $\varphi\in\mathcal{S}(\mathbb{R}^{2})$. Consequently, by Hölder’s inequality,
$\big|\mathbf{U}_{R}f\ast\varphi\big|^{p}\lesssim|\mathbf{U}_{R}f|^{p}\ast|\varphi|$
for $p\geq 1$. It follows that
$\int\big|\mathbf{U}_{R}f(x,t)\big|^{p}\,d\mu(x,t)\lesssim\int\big|\mathbf{U}_{R}f(x,t)\big|^{p}H(x,t)\,dxdt$
where $H=\mu\ast|\varphi|.$
We check that $H$ is $\alpha$-dimensional, using the dyadic decomposition $|\varphi|\lesssim\sum_{j\in\mathbb{N}}2^{-10j}1_{B_{2^{j}}}$. By using the decay and the assumption that $\langle\mu\rangle_{\alpha}\lesssim 1$, we have
$\int_{B_{\rho}(z)}H\lesssim\sum_{j\in\mathbb{N}}2^{-8j}\mu(B_{\rho}(z)+B_{2^{j}})\lesssim\rho^{\alpha}.$
Thus, $\langle H\rangle_{\alpha}\lesssim 1$. A similar computation shows that $\|H\|_{\infty}\lesssim 1$. Consequently, Theorem 2.2 and (5) give the desired estimate. ∎
[…]
(16) $\|e^{it\partial_{x}^{2}}f\|_{L^{p}(\mathbb{R}\times[0,1])}\leq C_{p,\gamma}\|f\|_{L^{p}_{\gamma}(\mathbb{R})},\;\;p\in(2,\infty)\;\text{and}\;\gamma>\max\Big(0,1-\frac{4}{p}\Big),$
which is due to Rogers [35]. Here $L^{p}_{\gamma}(\mathbb{R})$ denotes the $L^{p}$-Sobolev space equipped with the norm $\|f\|_{L_{\gamma}^{p}(\mathbb{R})}=\|(1-\Delta)^{\gamma/2}f\|_{L^{p}(\mathbb{R})}$. The regularity assumption on $\gamma$ in (16) is essentially sharp.
The bound obtained in 2.3 is essentially sharp when $p=2$ or $p=4$. Indeed,
(17) $\big\|\mathbf{U}_{R}f\big\|_{L^{p}(\mathbb{R}\times[0,R],\mu)}\lesssim R^{\zeta}\|f\|_{L^{p}(\mathbb{R})}$
holds only if
$\zeta\geq\begin{cases}\max\big(\frac{1}{2}-\frac{1}{p},~\frac{\alpha}{2p}\big),\;\;&\alpha\in[1,2],\\[3.0pt] \max\big(\frac{1}{2}-\frac{1}{p},~\frac{2\alpha-1}{2p}\big),\;\;&\alpha\in[0,1].\end{cases}$
We discuss the detail in Section 6.1.
2.4. Fractal local smoothing estimates relative to parabolic balls
The estimate (16) can be regarded as an analogue of the local smoothing phenomenon for the wave equation, first discovered by Sogge [37]. Indeed, comparing (16) with the sharp fixed-time estimate due to Miyachi [32],
$\|e^{it\partial_{x}^{2}}f\|_{L^{p}(\mathbb{R})}\lesssim_{p,\gamma}\|f\|_{L^{p}_{\gamma}(\mathbb{R})},\;\;p\in(1,\infty)\;\text{and}\;\gamma\geq\Big|1-\frac{2}{p}\Big|,$
it follows that averaging over a compact time interval yields a gain of $2/p$ derivatives whenever $p>4$. In the context of the Schrödinger equation, local smoothing estimates generally refer to such derivative gains obtained by averaging over a compact space-time region (see e.g. [36]).
Rogers [35] proved (16) by connecting it to the Fourier restriction estimate for the parabola. See [41] for a proof of (16) which relies on the square function estimate (1).
We seek to extend estimates of the form (16) to general measures on $\mathbb{R}\times[0,1]$ that satisfy suitable size conditions. We refer to these as fractal local smoothing estimates for the Schrödinger equation.
For $0\leq\beta\leq 3$, we consider a class of Borel measures on $\mathbb{R}^{2}$ for which
$[\mu]_{\beta,\text{par}}:=\sup_{z\in\mathbb{R}^{2},\rho>0}\rho^{-\beta}\mu\big(B_{\rho,\text{par}}(z)\big)\lesssim 1,$
where $B_{\rho,\textit{par}}(z)$ denotes the parabolic “ball” $(z_{1}-\rho,z_{1}+\rho)\times(z_{2}-\rho^{2},z_{2}+\rho^{2})$ for $z=(z_{1},z_{2})$. This class of measures naturally arises in view of the parabolic rescaling associated with the Schrödinger equation. For such measures, we consider the estimate
(18) $\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1],\mu)}\leq C[\mu]_{\beta,\text{par}}^{1/p}\|f\|_{L_{\gamma}^{p}(\mathbb{R})}.$
Theorem 2.4 (Parabolic $\beta$-dimensional case).
Let $0\leq\beta\leq 3$ and let $\mu$ be a Borel measure on $\mathbb{R}^{2}$ with $\mu]_{\beta,\mathrm{par}}\leq 1$.
  1. (i)
    (Sufficiency) For $2\leq p\leq 4$, there exists $C=C_{\beta,p,\gamma}>0$ such that (18) holds whenever
    $\gamma>\gamma_{\text{par}}(\beta):=\begin{cases}\frac{3-\beta}{4},&\beta\in[1,3],\\[1.72218pt] \frac{2-\beta}{2}-\frac{1-\beta}{p},&\beta\in[0,1].\end{cases}$
  2. (ii)
    (Necessity) Conversely, if (18) holds for some $0\leq\beta\leq 3$, then
    $\gamma\geq\begin{cases}\max\!\big\{1-\frac{\beta+1}{p},\,\frac{3-\beta}{2p}\big\},&\beta\in[1,3],\\[4.30554pt] \max\!\big\{1-\frac{\beta+1}{p},\,\frac{\beta}{p}\},&\beta\in[0,1].\end{cases}$
The necessary condition shows that $\gamma>\gamma_{\text{par}}(\beta)$ is essentially sharp for $p=4$ for all $0\leq\beta\leq 3$, and for $p=2$ for all $1\leq\beta\leq 3$. We prove the sufficiency part of Theorem 2.4 in Section 5 and the necessity part in Section 6.
2cm][c]0.46 $\tfrac{1}{p}$$\gamma$$\tfrac{1}{4}$$\tfrac{1}{2}$$\tfrac{2}{\beta+5}$
(a) $1\leq\beta\leq 3$
2cm][c]0.46 $\tfrac{1}{p}$$\gamma$$\tfrac{1}{4}$$\tfrac{1}{2}$$\tfrac{1}{2\beta+1}$
(b) $0\leq\beta\leq 1$
Figure 2. Sufficient (solid) and necessary (dotted) thresholds for Theorem 2.4.
We compare the $p=2$ case of Theorem 2.4 with known weighted Strichartz estimates. For the purpose, we consider the Morrey-Campanato type classes, which generalize the $L^{q}$ space. Given $\delta>0$ and $1\leq q\leq 3/\delta$, we define $\mathfrak{L}_{\text{par}}^{\delta,q}$ to be the set of nonnegative weights $H\in L_{\text{loc}}^{q}(\mathbb{R}\times\mathbb{R})$, equipped with the norm
$\|H\|_{\mathfrak{L}_{\text{par}}^{\delta,q}}:=\sup_{(x,t)\in\mathbb{R}^{1+1},r>0}r^{\delta}\Big(\frac{1}{r^{3}}\int_{B_{r,\text{par}}(x,t)}H(y,s)^{q}\,dyds\Big)^{1/q}.$
For instance, $|(x,t)|^{-3/q}\in\mathfrak{L}_{\text{par}}^{\delta,q}$ for $q<3/\delta$, although it does not belong to $L^{q}$ space. In fact, $L^{q}=\mathfrak{L}_{\text{par}}^{\delta,q}$ when $\delta=3/q$, and $L^{q,\infty}\subset\mathfrak{L}_{\text{par}}^{\delta,q}$ when $\delta<3/q$. Moreover, for $\mu=Hdxdt$, we have
$\mu]_{\beta,\text{par}}=\|H\|_{\mathfrak{L}_{\text{par}}^{3-\beta,1}}.$
Barceló et al. [3] established weighted Strichartz estimates of the form
(19) $\|e^{it\partial_{x}^{2}}f\|_{L_{x,t}^{2}(H(x,t))}\leq C\|H\|_{\mathfrak{L}_{\text{par}}^{2\gamma+2,q}}^{1/2}\|f\|_{\dot{H}^{\gamma}}$
where the exponent $2\gamma+2$ is determined by the scaling invariance. In [3], it was shown that (19) holds for $\frac{1}{4}\leq\gamma<\frac{1}{2}$ and $1<q\leq\frac{3}{2\gamma+2}$ (with higher dimensional analogue). See [3, 27] for the case $0\leq\gamma<\frac{1}{4}$ and [2] for results with time-dependent weights $H$.
Note that when $0<\beta\leq 1/2$, (19) yields
(20) $\|e^{it\partial_{x}^{2}}f\|_{L_{x,t}^{2}(H(x,t))}\leq C\|H\|_{\mathfrak{L}_{\text{par}}^{3-\beta,q}}^{1/2}\|f\|_{\dot{H}^{\gamma}}$
for $\gamma=(1-\beta)/2$ and $1<q\leq\frac{3}{3-\beta}$. Since $q>1$, (20) does not seem to imply estimates (18) for measures $\mu=Hdxdt$. Nevertheless, (20) is superior to Theorem 2.4 for the range $0<\beta\leq 1/2$ in the sense that it provides a global estimate in both space and time and the regularity index $\gamma$ is optimal, matching the necessary condition in Theorem 2.4 $(ii)$.
2.5. Fractal local smoothing estimates
We now consider fractal local smoothing estimates for the Schrödinger operator with $\alpha$-dimensional measures in the standard (non-parabolic) sense. Let $0\leq\alpha\leq 2$ and consider a class of Borel measures on $\mathbb{R}^{2}$ for which
(21) $[\mu]_{\alpha}:=\sup_{z\in\mathbb{R}^{2},\;\rho>0}\rho^{-\alpha}\mu\big(B_{\rho}(z)\big)<\infty.$
Here, the condition (21) differs slightly from $\langle\mu\rangle_{\alpha}$ defined in (15), in that the supremum is taken over all $\rho>0$, rather than $\rho>1$.
In the $L^{2}$ setting, a variety of results for general fractal measures are known (see, e.g., [40, 31, 34, 12], but much less is understood beyond the $L^{2}$ framework. For product measures, however, optimal results have been obtained by Lee–Lee–Roncal [28], with further developments connected to Assouad dimensions. Suppose $\nu$ is supported on $[0,1]$ and satisfies
$\nu((t-\rho,t+\rho))\leq C\rho^{\alpha},\quad(t,\rho)\in\mathbb{R}\times\mathbb{R}_{+}.$
In [28, Theorem 1.6], it is shown that if $\gamma\geq-\frac{\alpha}{2}$, then $H^{\gamma}$$L_{t}^{2}(d\nu;L_{x}^{2}(-1,1))$ estimates holds for $e^{it\partial_{x}^{2}}f$. Moreover, optimal results in higher dimensions and weighted Strichartz-type estimates of the form $H^{\gamma}$$L_{t}^{q}(d\nu;L_{x}^{r})$ for fractional Schrödinger operators are established in [28] (see also [4] for analogous results for the wave operator).
For the wave equation, the situation is better understood: not only are the product-type estimates optimal (see [4]), but there are also extensive results on $L^{p}-L^{q}$ estimates with respect to more general fractal measures (see [34, 8, 25, 23, 7]; see also [24, 22] for related results in the case of product measures).
We consider the local smoothing type estimates for the Schrödinger operator relative to fractal measure $\mu$:
(22) $\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1],\mu)}\leq C[\mu]_{\alpha}^{1/p}\|f\|_{L_{\gamma}^{p}(\mathbb{R})}$
for some $C=C_{\alpha,p,\gamma}$. By combining the weighted square function estimates in Theorem 1.1 with the standard strategy, we obtain the following.
Theorem 2.5.
Let $0\leq\alpha\leq 2$, and let $\mu$ be a Borel measure on $\mathbb{R}^{2}$ such that $\mu]_{\alpha}\leq 1$.
  1. (i)
    (Sufficiency) For $2\leq p\leq 4$, the estimate (22) holds whenever
    $\gamma>\gamma(\alpha,p):=\begin{cases}\frac{2-\alpha}{2},&\alpha\in[1,2],\\[3.44444pt] \frac{2-\alpha}{2}+\frac{\alpha-1}{p},&\alpha\in[0,1].\end{cases}$
  2. (ii)
    (Necessity) Conversely, (22) can hold only if
    $\gamma\geq\begin{cases}\max\!\big(1-\frac{2\alpha}{p},\,\frac{2-\alpha}{p}\big),&\alpha\in[1,2],\\[3.44444pt] \max\!\big(1-\frac{\alpha+1}{p},\,\frac{\alpha}{p}\big),&\alpha\in[0,1].\end{cases}$
The sufficient conditions $\gamma(\alpha,p)$ are essentially sharp for $p=4$ for all $\alpha\in[0,2]$ and $p=2$ for $\alpha\in[1,2]$. We prove the sufficiency part of Theorem 2.4 in Section 5 and the necessity part in Section 6.
3. More examples
3.1. A lower bound for the weighted square function estimate
We give a lower bound for the weighted square function estimate by using an example from [1] discussed in Remark 1.3.
We fix a parameter $0<\kappa\leq 1/2$. For each $l\in R^{-\kappa}\mathbb{Z}\cap[-1/2,1/2]$, let $\Omega_{l}=[l-R^{-1},l+R^{-1}]$ and
$f_{l}(x,t)=\eta(R^{-1}t)\eta(R^{-1}x)R\int e^{i(x\xi+t\xi^{2})}1_{\Omega_{l}}(\xi)d\xi,$
where $\eta$ is defined in Section 2.3. It follows that $|f_{l}(x,t)|\sim 1$ on $B_{R}$ and decays rapidly away from $B_{R}$. Thus,
$\big\|\big(\sum_{l}|f_{l}|^{2}\big)^{1/2}\big\|_{L^{p}}\lesssim R^{\kappa/2}R^{2/p}.$
Let $\Omega=\cup_{l}\Omega_{l}$ and $f=\sum_{l}f_{l}$. For a sufficiently small $0<c<1$, define
(23) $\Gamma=(2\pi R^{\kappa}\mathbb{Z}\times 2\pi R^{2\kappa}\mathbb{Z})\cap B_{cR}(0),\;\;Y=\Gamma+B_{c}(0).$
One can check that $|Y\cap B_{\rho}|\lesssim\rho^{2-3\kappa}$ for all $\rho\geq 1$. Therefore, if we let $\alpha=2-3\kappa$, then $H=1_{Y}$ is an $\alpha$-dimensional weight. Moreover, we have $x\xi+t\xi^{2}\in 2\pi\mathbb{Z}+B_{0.01}(0)$ whenever $(x,t)\in Y$ and $\xi\in\Omega$. Consequently,
$\|f\|_{L^{p}(Y)}\sim R^{\kappa}|Y|^{1/p}\sim R^{\kappa+(2-3\kappa)/p}.$
Combining these estimates, we get the lower bound
(24) $\|f\|_{L^{p}(Y)}/\big\|\big(\sum_{l}|f_{l}|^{2}\big)^{1/2}\big\|_{L^{p}}\gtrsim R^{\kappa(\frac{1}{2}-\frac{3}{p})}=R^{-(2-\alpha)(\frac{1}{p}-\frac{1}{6})}.$
3.2. A sharp example beyond interpolation
In view of 2.3 concerned with $\alpha$-dimensional weights or measures, it seems natural to ask whether our weighted $L^{p}$-estimates, Theorem 2.1 and Theorem 2.2 expressed in terms of $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)$, can yield results beyond what can be obtained by interpolating between the $L^{2}$ and $L^{4}$ estimates that they provide. The following example shows that the answer is affirmative. For this particular weight, the dominant scale $R^{-1/2}\leq s\leq 1$ for $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)$ depends on the exponent $p$, being either 1 or $R^{-1/2}$.
Example 4.
Let $1<\alpha<2$ and $p_{\alpha}=4/(3-\alpha)$. We construct a positive weight $H=1_{Y}$ in $\mathbb{R}^{2}$ such that $\langle H\rangle_{\alpha}\lesssim 1$ and
(25) $\max_{U\in\mathbb{U}}\kappa_{p,H}(U)\lesssim\begin{cases}R^{-\frac{2-\alpha}{2p}},&2\leq p\leq p_{\alpha}\\[4.30554pt] R^{-\frac{3-\alpha}{2}(\frac{1}{p}-\frac{1}{4})},&p_{\alpha}\leq p\leq 4.\end{cases}$
Note that for the weight in 4, Theorem 2.1 and Theorem 2.2 yield $L^{p}$-estimates which cannot be obtained by interpolating between $L^{2}$ and $L^{4}$ estimates for $2<p<4$.
Regarding Theorem 1.1, let $\sigma$ denote the infimum of exponents for which the following estimate holds with the specific weight $H=1_{Y}$ to be defined:
$\|f\|_{L^{p}(Hdx)}\lessapprox R^{\sigma}\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{\frac{1}{2}}\big\|_{L^{p}(\mathbb{R}^{2})}.$
Applying (25) to Theorem 1.1 yields that
(26) $\sigma\leq\max\Big\{-\frac{2-\alpha}{2p},~-\frac{3-\alpha}{2}\Big(\frac{1}{p}-\frac{1}{4}\Big)\Big\}.$
Note that the upper bound (26) is strictly stronger than the one
(27) $\sigma\leq-(2-\alpha)\Big(\frac{1}{p}-\frac{1}{4}\Big)$
given by (5) for all intermediate $2<p<4$.
The upper bound (26) is sharp for $2\leq p\leq p_{\alpha}$ and $p=4$. This can be seen from the lower bound
(28) $\sigma\geq\max\Big\{-\frac{2-\alpha}{2p},~-2\big(\frac{1}{p}-\frac{1}{4}\big)\Big\}.$
The lower bound follows from examples in Remark 1.3. Specifically, the unit ball example provides the lower bound $\sigma\geq-2\big(\frac{1}{p}-\frac{1}{4}\big)$, while the single wave packet example yields the lower bound $\sigma\geq-\frac{2-\alpha}{2p}$ since $Y\subset T$ for a single $R^{1/2}\times R$ tube $T$ with $|Y|/|T|\sim R^{-(2-\alpha)/2}$. On the other hand, the computation from Section 3.1 yields a lower bound weaker than (28).
We compare lower and upper bounds for $\sigma$ in Figure 3.
$\sigma$$\tfrac{1}{p}$$\tfrac{1}{4}$$\tfrac{1}{2}$$\tfrac{1}{\alpha+2}$$\tfrac{3-\alpha}{4}$(27)(28) (26)(28)
Figure 3. Blue and black lines denote the upper bounds on from (26) and (27), respectively, while the dotted line indicates the lower bound from (28).
A weight $H$ satisfying (25).
Let $\kappa\in(0,\frac{1}{6})$ and $\alpha:=2-6\kappa\in(1,2)$. Let
$Y=\Gamma\cap([0,R^{1/2}]\times[0,R])+B_{c}(0),$
where $\Gamma$ is defined in (23). Let $H=1_{Y}$. For any $z\in\mathbb{R}^{2}$ and $\rho\geq 1$, we have
$H(B_{\rho}(z))=|Y\cap B_{\rho}(z)|$
$\lesssim\begin{cases}1,&\rho\in[1,R^{\kappa}],\\[4.30554pt] \rho/R^{\kappa},&\rho\in[R^{\kappa},R^{2\kappa}],\\[4.30554pt] \rho^{2}/R^{3\kappa},&\rho\in[R^{2\kappa},R^{1/2}],\end{cases}\qquad\begin{cases}(R^{1/2}/R^{\kappa})(\rho/R^{2\kappa}),&\rho\in[R^{1/2},R],\\[4.30554pt] R^{3/2}/R^{3\kappa},&\rho\in[R,\infty).\end{cases}$
It follows that $|Y\cap B_{\rho}|\lesssim\rho^{2-6\kappa}$ for all $\rho\geq 1$. Thus $H$ is $\alpha=(2-6\kappa)$-dimensional.
Next, we prove (25) by computing $H(T)$ and $H(U)$ directly. Let $T\in\mathbb{T}_{\tau}$ and $U\in\mathbb{U}_{\tau}$ for some $|\tau|=s$. We have
$\frac{H(U)}{|U|}=\frac{|U\cap Y|}{|U|}\leq\frac{|Y|}{|U|}\sim\frac{R^{\frac{3}{2}-3\kappa}}{R^{2}s}.$
Regarding $T$, we have
$\frac{H(T)}{|T|}=\frac{|T\cap Y|}{|T|}\lesssim\begin{cases}s^{-3}R^{-3\kappa}/s^{-3},&s\in[R^{-1/2},R^{-\kappa}],\\ 1/s^{-3},&s\in[R^{-\kappa},1].\end{cases}$
Therefore,
$\max_{U\in\mathbb{U}}\kappa_{p,H}(U)$ $\lesssim\max\Big(\max_{R^{-\frac{1}{2}}\leq s\leq R^{-\kappa}}R^{-\frac{3\kappa}{4}}(s^{-1}R^{-(\frac{1}{2}+3\kappa)})^{\frac{1}{p}-\frac{1}{4}},\max_{R^{-\kappa}\leq s\leq 1}s^{\frac{3}{4}}(s^{-1}R^{-(\frac{1}{2}+3\kappa)})^{\frac{1}{p}-\frac{1}{4}}\Big).$
Recall that $2\leq p\leq 4$. Substituting $\kappa=(2-\alpha)/6$ yields
$\max_{U\in\mathbb{U}}\kappa_{p,H}(U)$ $\lesssim\max\Big(R^{-\frac{3\kappa}{p}},R^{-(\frac{1}{2}+3\kappa)(\frac{1}{p}-\frac{1}{4})}\Big)=\max\Big(R^{-\frac{2-\alpha}{2p}},~R^{-\frac{3-\alpha}{2}(\frac{1}{p}-\frac{1}{4})}\Big),$
which is equivalent to (25). ∎
4. Weighted wave envelope estimates
In this section, we prove Theorem 1.4 via bilinear restriction theorem.
4.1. Reductions
Let $H$ be a positive weight. We first reduce Theorem 1.4 to the special case $H=1_{Y}$ for a subset $Y\subset\mathbb{R}^{2}$, namely,
(29) $\|f\|_{L^{p}(Y)}^{p}\lessapprox\sum_{R^{-1/2}\leq s\leq 1}\sum_{|\tau|=s}\sum_{U\in\mathbb{U}_{\tau}}\kappa_{p,Y}(U)^{p}|U|^{1-\frac{p}{2}}\|(\sum_{\theta\subset\tau}|f_{\theta}|^{2})^{1/2}\|_{L^{2}(w_{U})}^{p}.$
Here and in the following, we use the notation $\kappa_{p,Y}(U)$ for $\kappa_{p,1_{Y}}(U)$ when $Y\subset\mathbb{R}^{2}$.
Assume that (29) holds. By the assumption on $H$ and a dyadic pigeonholing, there exists a dyadic number $\lambda\in[R^{-400},1]$ and a subset $Y_{\lambda}\subset\mathbb{R}^{2}$ such that $H(x)\sim\lambda$ for $x\in Y_{\lambda}$ and
(30) $\|f\|_{L^{p}(Hdx)}^{p}\leq C(\log R)^{O(1)}\lambda\|f\|_{L^{p}(Y_{\lambda})}^{p}.$
Moreover, for any $E\subset\mathbb{R}^{2}$, we have $\lambda|E\cap Y_{\lambda}|\sim H(E\cap Y_{\lambda})\leq H(E)$. Hence, by the definition of (3),
(31) $\lambda\,\kappa_{p,{Y_{\lambda}}}(U)^{p}\lesssim\kappa_{p,H}(U)^{p}.$
Using (30) and (31), we conclude that (29) implies (8). Therefore, the proof of Theorem 1.4 reduces to establishing (29).
For the rest of the section, we prove (29).
4.2. Decomposition
For $R\geq 1$, let $K$ be a dyadic number such that $K\sim\log R$. We choose $m\in\mathbb{N}$ such that $K^{m}\sim R^{1/2}$. We consider larger canonical covering $\tau$ such that $|\tau|=s$ for each scale
$s=K^{-j},\quad j=0,1,\cdots,m.$
At the smallest scale for $j=m$, we obtain the collection of blocks $\{\theta\}$ of size $|\theta|=K^{-m}\sim R^{-1/2}.$ For each block $\tau$ at an intermediate scale, we define
$f_{\tau}=\sum_{\theta\subset\tau}f_{\theta}.$
We use the following pointwise estimate from [10, Section 5].
Lemma 4.1.
For any $x\in\mathbb{R}^{2}$, there is an absolute constant $C>0$ (independent of $x$) such that
(32) $|f(x)|^{p}\leq C^{m}\sum_{\theta}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{R^{-1/2}<s\leq 1}\sum_{|\tau|=s}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-1}s\\ d(\tau_{1},\tau_{2})\geq\frac{1}{2}K^{-1}s\end{subarray}}|f_{\tau_{1}}(x)f_{\tau_{2}}(x)|^{\frac{p}{2}},$
where $d(\tau_{1},\tau_{2})$ denotes the distance between $\tau_{1}$ and $\tau_{2}$.
The inequality follows from the Bourgain–Guth argument [5]. For the convenience of the reader, we provide a proof in Appendix A.
By integrating (32) over $Y$, we get
(33) $\|f\|_{L^{p}(Y)}^{p}\!\leq\!C^{m}\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}+C^{m}K^{p}\sum_{R^{-1/2}<s\leq 1}\sum_{|\tau|=s}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-1}s\\ d(\tau_{1},\tau_{2})\geq\frac{1}{2}K^{-1}s\end{subarray}}\!\!\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p}.$
By the computation (9), the first term on the right-hand side of (33) is bounded by
$\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}\lesssim\sum_{\theta}\sum_{U\in\mathbb{U}_{\theta}}\frac{|U\cap Y|}{|U|}|U|^{1-\frac{p}{2}}\|f_{\theta}\|_{L^{2}(w_{U})}^{p}.$
This corresponds to the contribution at the smallest scale $s=R^{-1/2}$ in (8) (cf. (10)).
Thus it remains to control the bilinear terms, which reduces to the following proposition.
Proposition 4.2.
Let $1\leq K\ll R$ and $\tau_{1},\tau_{2}\subset\tau$ satisfy $|\tau_{j}|=K^{-1}|\tau|$, $j=1,2$, and $d(\tau_{1},\tau_{2})\geq K^{-1}|\tau|$. Then for any $\epsilon>0$, $2\leq p\leq 4$ and $U\in\mathbb{U}_{\tau}$, we have
(34) $\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(U\cap Y)}^{p}\lessapprox\,\kappa_{p,Y}(U)^{p}|U|^{1-\frac{p}{2}}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{2}(w_{U})}^{p}.$
Summing (34) over all $U\subset\mathbb{U}_{\tau}$ yields
$\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p}\lessapprox\sum_{U\in\mathbb{U}_{\tau}}\kappa_{p,Y}(U)^{p}|U|^{1-\frac{p}{2}}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{2}(w_{U})}^{p}.$
Since for each fixed $\tau$ there are only $K^{O(1)}$ (and hence $\lesssim\log R$) such pairs $\tau_{1},\tau_{2}\subset\tau$, this logarithmic factor can be absorbed into the $O((\log R)^{O(1)})$ loss. Consequently, the bilinear term in the sum in (33) is bounded by the right-hand side of (8).
4.3. Proof of 4.2
We first reduce the matter to treat the case $p=4$ for (34), which can be handled by the bilinear restriction theorem. We claim that 4.2 holds with $p=4$:
(35) $\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{4}(U\cap Y)}\lessapprox\kappa_{4,Y}(U)|U|^{-\frac{1}{4}}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{2}(w_{U})}.$
Here, by definition (3), when $p=4$ we have
(36) $\kappa_{4,Y}(U)^{4}=\max_{\begin{subarray}{c}T\in\mathbb{T}_{\tau(U)}:\\ T\subset U\end{subarray}}\frac{|T\cap Y|}{|T|}.$
Having established 4.2 for $p=4$, it remains to treat the range $2\leq p\leq 4$. By Hölder’s inequality,
(37) $\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(U\cap Y)}\leq|U\cap Y|^{\frac{1}{p}-\frac{1}{4}}\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{4}(U\cap Y)}.$
We combine (35) and (37), and observe that
$\kappa_{4,Y}(U)|U|^{-\frac{1}{4}}|U\cap Y|^{\frac{1}{p}-\frac{1}{4}}=\max_{\begin{subarray}{c}T\in\mathbb{T}_{\tau(U)}:\\ T\subset U\end{subarray}}\Big(\frac{|T\cap Y|}{|T|}\Big)^{\frac{1}{4}}\Big(\frac{|U\cap Y|}{|U|}\Big)^{\frac{1}{p}-\frac{1}{4}}|U|^{\frac{1}{p}-\frac{1}{2}}.$
This yields the desired estimate (34), thereby completing the proof of Theorem 1.4. ∎
4.4. Bilinear restriction estimates
It remains to establish (35). To this end, we apply parabolic rescaling. Fix $\tau$ with $|\tau|=s$ centered at $(c,c^{2})$. Let $A_{\tau}$ denote the affine transform
$A_{\tau}(\xi_{1},\xi_{2})=(c,c^{2})+(s\xi_{1},2cs\xi_{1}+s^{2}\xi_{2}).$
Thus $\tau$ can be identified with the image of $1,1]\times[-2,2]$ under $A_{\tau}$. For each $\theta\subset\tau$, we define $g_{\theta}$ by
$\widehat{g_{\theta}}(\xi)=s^{3}\widehat{f_{\theta}}(A_{\tau}\xi)$
and let
$g_{\tau_{i}}:=\sum_{\theta\subset\tau_{i}}g_{\theta},\quad i=1,2.$
With the new scale $R_{s}:=Rs^{2}$, $\widehat{g_{\theta}}$ is supported on a canonical box of dimensions $R_{s}^{-1/2}\times R_{s}^{-1}$ covering the $R_{s}^{-1}$-neighborhood of the parabola $\mathcal{P}$. Moreover, the supports of $\widehat{g_{\tau_{1}}},\widehat{g_{\tau_{2}}}$ are $K^{-1}$-separated.
We may write $g_{\theta}(x)=c_{\tau}(x)f_{\theta}(L_{\tau}x)$, where $|c_{\tau}(x)|=1$ and $L_{\tau}$ is the linear transform
(38) $L_{\tau}=\begin{pmatrix}s^{-1}&-2cs^{-2}\\ 0&s^{-2}\end{pmatrix}$.
Let $B=L_{\tau}^{-1}(U)$, which is a cube of side length $R_{s}$.
For any unit cube $q\subset B$ (so $|q|=1$), recalling the definition (2) and applying the change of variables $x\rightarrow L_{\tau}x$, the quantity $\kappa_{4,Y}(U)$ given in (36) can be written as
$\kappa_{4,Y}(U)^{4}=\max_{q\subset B}\frac{|L_{\tau}(q)\cap Y|}{|L_{\tau}(q)|}=\max_{q\subset B}\frac{|q\cap\widetilde{Y}|}{|q|}$
where $\widetilde{Y}=L_{\tau}^{-1}(Y)$. With this rescaling, and using the identity $|B|^{-1}=|U|^{-1}|\det L_{\tau}|$, the estimate (35) reduces to prove the following.
Lemma 4.3.
Let $B$ and $g_{\tau_{i}}$ be as above. Let $q$ be a unit cube contained in $B$. Then
(39) $\int_{B\cap\widetilde{Y}}|g_{\tau_{1}}g_{\tau_{2}}|^{2}\,\lessapprox\max_{q\subset B}\frac{|q\cap\widetilde{Y}|}{|q|}|B|^{-1}\|g_{\tau_{1}}\|_{L^{2}(w_{B})}^{2}\|g_{\tau_{2}}\|_{L^{2}(w_{B})}^{2}$
where $w_{B}$ is an $L^{\infty}$-normalized weight which decays rapidly away from $B$.
By the local $L^{2}$-orthogonality, for $\tau_{1},\tau_{2}\subset\tau$ we have
(40) $\|g_{\tau_{i}}\|_{L^{2}(w_{B})}\lesssim\big\|\big(\sum_{\theta\subset\tau}|g_{\theta}|^{2}\big)^{1/2}\big\|_{L^{2}(w_{B})},\quad i=1,2,$
Thus, combining (39) with (40), we obtain (35).
It remains to prove Lemma 4.3. For this purpose, we invoke the following local bilinear restriction estimate for the parabola:
Theorem 4.4.
Let $1\leq K\ll r$ and $B$ be a ball of radius $r$. Let $\tau_{1}$ and $\tau_{2}$ be boxes contained in $N_{r^{-1}}(\mathcal{P})$ such that $d(\tau_{1},\tau_{2})\geq K^{-1}$. If $g_{\tau_{1}}$ and $g_{\tau_{2}}$ are Fourier supported on $\tau_{1}$ and $\tau_{2}$, respectively, then
$\int_{B}|g_{\tau_{1}}g_{\tau_{2}}|^{2}\lesssim K^{O(1)}|B|^{-1}\|g_{\tau_{1}}\|_{L^{2}(w_{B})}^{2}\|g_{\tau_{2}}\|_{L^{2}(w_{B})}^{2},$
where $w_{B}$ is a $L^{\infty}$-normalized weight which decays rapidly away from $B$.
The proof of Theorem 4.4 is standard; see, for example, [38] or [29, Lemma 2.4].
We now turn to the proof of Lemma 4.3.
Proof of Lemma 4.3.
Since $|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}$ has compact Fourier support, it is locally constant on unit cubes. More precisely, we may choose $\phi=\phi_{N}:=(1+|\cdot|)^{-N}$ for some sufficiently large $N\in\mathbb{N}$ such that
$|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}\lesssim|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}*\phi.$
Note that the convolution $|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}*\phi$ is locally constant on unit cubes in the sense that $|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}*\phi(x)\sim|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}*\phi(y)$ whenever $|x-y|\lesssim 1$. Hence, for each unit cube $q\subset B$,
$\int_{q\cap\widetilde{Y}}|g_{\tau_{1}}g_{\tau_{2}}|^{2}\lesssim\frac{|q\cap\widetilde{Y}|}{|q|}\int_{q}|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}*\phi(x)dx.$
Summing this inequality over $q\subset B$ yields
(41) $\int_{B\cap\widetilde{Y}}|g_{\tau_{1}}g_{\tau_{2}}|^{2}\lesssim\max_{q\subset B}\frac{|q\cap\widetilde{Y}|}{|q|}\int_{B}|g_{\tau_{1}}|^{2}|g_{\tau_{2}}|^{2}*\phi(x)dx.$
The integral on the right-hand side of (41) can be written as
$\int|g_{\tau_{1}}g_{\tau_{2}}|^{2}\phi*1_{B},$
where $\phi*1_{B}$ is a $L^{\infty}$-normalized weight which decays rapidly away from $B$. Using this decay property and applying Theorem 4.4 to (41), we get the desired bound in Lemma 4.3. ∎
5. Fractal local smoothing estimates
In this section, we establish the sufficiency parts in Theorem 2.4 and Theorem 2.5.
5.1. Rescaling and reductions
We reduce Theorem 2.4 and Theorem 2.5 to 2.3. By homogeneity, we may assume $\mu]_{\alpha}=1$ and $\mu]_{\beta,\text{par}}=1$ for the proofs of Theorem 2.4 and Theorem 2.5, respectively.
Fix $2\leq p\leq 4$. Given a measure $\mu$, we need to verify
$\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1];\mu)}\lesssim\|f\|_{L^{p}_{\gamma}(\mathbb{R})}$
for all $\gamma>\gamma_{\text{par}}(\beta)$ and $\gamma>\gamma(\alpha,p)$ respectively. By a standard Littlewood–Paley reduction, it suffices to show that
$\|e^{it\partial_{x}^{2}}f\|_{L^{p}(\mathbb{R}\times[0,1];\mu)}\lessapprox R^{\gamma}\|f\|_{p}$
for any function $f$ whose Fourier transform is supported on $\{\xi\in\mathbb{R}:|\xi|\leq R\}$.
For the purpose, we define $f_{R}(x)=f(R^{-1}x)$ so that $\widehat{f_{R}}$ is supported on [1,1][-1,1] and $\|f_{R}\|_{p}=R^{\frac{1}{p}}\|f\|_{p}$. A change of variable $\xi\rightarrow R\xi$ gives
$|e^{it\partial_{x}^{2}}f(x)|\sim|\mathbf{U}_{R^{2}}f_{R}(Rx,R^{2}t)|,\;\;(x,t)\in\mathbb{R}\times[0,1],$
where $\mathbf{U}_{R^{2}}$ is defined in (13). We also define the rescaled measure $\mu_{R}$ on $\mathbb{R}^{2}$ by $\mu_{R}(E)=\int 1_{E}(Rx,R^{2}t)d\mu(x,t)$ so that we have
$\int|\mathbf{U}_{R^{2}}f_{R}|^{p}(x,t)\,d\mu_{R}(x,t)=\int|\mathbf{U}_{R^{2}}f_{R}|^{p}(Rx,R^{2}t)\,d\mu(x,t).$
Thus, we have
(42) $\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1];\mu)}\lesssim\big\|\mathbf{U}_{R^{2}}f_{R}\big\|_{L^{p}(\mathbb{R}^{2};\mu_{R})}.$
Lemma 5.1.
For $\mu_{R}$ defined as above, we have
  1. (1)
    If $\mu]_{\beta,\text{par}}\leq 1$, then
    $\langle\mu_{R}\rangle_{(\beta+1)/2}$ $\lesssim R^{-\beta},\;\;\;\beta\in[1,3],$
    $\langle\mu_{R}\rangle_{\beta}$ $\lesssim R^{-\beta},\;\;\;\beta\in[0,1].$
  2. (2)
    If $\mu]_{\alpha}\leq 1$, then
    $\langle\mu_{R}\rangle_{\alpha}\lesssim\begin{cases}R^{1-2\alpha},&\alpha\in[1,2],\\ R^{-\alpha},&\alpha\in[0,1].\end{cases}.$
Proof.
Consider the ball $B_{\rho}(z)$ of radius $\rho\geq 1$ for $z=(z_{1},z_{2})$. By the definition of $\mu_{R}$, we have
$\mu_{R}\big(B_{\rho}(z)\big)\leq\mu\Big(\big(\frac{z_{1}-\rho}{R},\frac{z_{1}+\rho}{R}\big)\times\big(\frac{z_{2}-\rho}{R^{2}},\frac{z_{2}+\rho}{R^{2}}\big)\Big).$
Suppose that $\mu]_{\beta,\text{par}}=1$. We may cover the rectangle $\big(\frac{z_{1}-\rho}{R},\frac{z_{1}+\rho}{R}\big)\times\big(\frac{z_{2}-\rho}{R^{2}},\frac{z_{2}+\rho}{R^{2}}\big)$ by $O(\rho^{1/2})$ parabolic rectangles of dimensions $\rho^{1/2}/R\times\rho/R^{2}$. Alternatively, we can just cover it by a parabolic rectangle of dimensions $\rho/R\times(\rho/R)^{2}$. Therefore, we have
$\mu_{R}\big(B_{\rho}(z)\big)\lesssim\min\big(\rho^{\frac{1}{2}}\big(\rho^{\frac{1}{2}}/R\big)^{\beta},(\rho/R)^{\beta}\big)=R^{-\beta}\min(\rho^{\frac{1+\beta}{2}},\rho^{\beta}),$
which verifies the claim.
Next, assume that $\mu]_{\alpha}=1$. We may cover the rectangle $\big(\frac{z_{1}-\rho}{R},\frac{z_{1}+\rho}{R}\big)\times\big(\frac{z_{2}-\rho}{R^{2}},\frac{z_{2}+\rho}{R^{2}}\big)$ by a single ball of radius $\sim\rho/R$, or alternatively, $O(R)$ balls of radius $\rho/R^{2}$. Thus,
$\mu_{R}\big(B_{\rho}(z)\big)\lesssim\min\big(R\big(\rho/R^{2}\big)^{\alpha},\big(\rho/R\big)^{\alpha}\big)=\min\big(R^{1-2\alpha},R^{-\alpha}\big)\rho^{\alpha}.\qed$
5.2. Proof of sufficient conditions in Theorem 2.4 and Theorem 2.5
We prove sufficient conditions in Theorem 2.4 and Theorem 2.5. Let $f$ be a function whose Fourier transform is supported on [R,R][-R,R], $R\geq 1$. By the reduction in (42), we have
(43) $\|e^{it\partial_{x}^{2}}f\|_{L^{p}(\mathbb{R}\times[0,1];\mu)}\lesssim R^{-\frac{\theta}{p}}\|\mathbf{U}_{R^{2}}f_{R}\|_{L^{p}(\mathbb{R}^{2},R^{\theta}\mu_{R})}$
where $\theta\in\mathbb{R}$ is to be chosen depending on the measure $\mu$.
5.2.1. Proof of Theorem 2.4
Suppose that $\mu]_{\beta,\text{par}}=1$. By Lemma 5.1, we have $\langle R^{\beta}\mu_{R}\rangle_{(\beta+1)/2}\lesssim 1$ when $\beta\in[1,3]$ and $\langle R^{\beta}\mu_{R}\rangle_{\beta}\lesssim 1$ when $\beta\in[0,1]$. Thus, 2.3 yields
$\big\|\mathbf{U}_{R^{2}}f_{R}\big\|_{L^{p}(\mathbb{R}^{2},R^{\beta}\mu_{R})}\lessapprox\begin{cases}R^{2\left(\frac{1}{p}-(2-\frac{\beta+1}{2})\left(\frac{1}{p}-\frac{1}{4}\right)\right)}\|f_{R}\|_{L^{p}}=R^{\frac{3-\beta}{4}+\frac{\beta-1}{p}}\|f_{R}\|_{L^{p}},&\quad\beta\in[1,3],\\ R^{2\left(\frac{1}{p}-(2-\beta)\left(\frac{1}{p}-\frac{1}{4}\right)\right)}\|f_{R}\|_{L^{p}}=R^{\frac{2-\beta}{2}-\frac{2(1-\beta)}{p}}\|f_{R}\|_{L^{p}},&\quad\beta\in[0,1].\end{cases}$
Recalling (43) with $\theta=\beta$ and $\|f_{R}\|_{p}=R^{1/p}\|f\|_{p}$, we get
$\|e^{it\partial_{x}^{2}}f\|_{L^{p}(\mathbb{R}\times[0,1];\mu)}\lessapprox\begin{cases}R^{\frac{3-\beta}{4}}\|f\|_{L^{p}},\quad&\beta\in[1,3],\\[3.44444pt] R^{\frac{2-\beta}{2}-\frac{1-\beta}{p}}\|f\|_{L^{p}},\quad&\beta\in[0,1].\end{cases}$
This completes the proof of the sufficiency part in Theorem 2.4. ∎
5.2.2. Proof of Theorem 2.5
The proof the sufficient condition in Theorem 2.5 for $\alpha\in[0,1]$ is the same as the proof for the case $\beta\in[0,1]$ of Theorem 2.4. For $\alpha\in[1,2]$, by Lemma 5.1, we have $\langle R^{2\alpha-1}\mu_{R}\rangle_{\alpha}\lesssim 1$. Thus, 2.3 yields
$\big\|\mathbf{U}_{R^{2}}f_{R}\big\|_{L^{p}(\mathbb{R}^{2},R^{2\alpha-1}\mu_{R})}\lessapprox R^{2\left(\frac{1}{p}-(2-\alpha)\left(\frac{1}{p}-\frac{1}{4}\right)\right)}\|f_{R}\|_{L^{p}}.$
Consequently, by (43) we have
$\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1];\mu)}\lesssim R^{-\frac{2\alpha-1}{p}}\|\mathbf{U}_{R^{2}}f_{R}\|_{L^{p}(\mathbb{R}^{2},R^{2\alpha-1}\mu_{R})}\lessapprox R^{\frac{2-\alpha}{2}-\frac{1}{p}}\|f_{R}\|_{L^{p}}=R^{\frac{2-\alpha}{2}}\|f\|_{L^{p}}.$
This completes the proof. ∎
5.3. A sketch of the proof of (14)
We restate (14) with replacing $R$ by $R^{2}$:
Proposition 5.2.
For $2\leq p\leq 4$, we have
(44) $\Big\|\big(\sum_{J}|\mathbf{U}_{R^{2}}f_{J}|^{2}\big)^{1/2}\Big\|_{L^{p}}\lessapprox R^{\frac{2}{p}}\|f\|_{L^{p}}.$
A proof of 5.2 is given in [41]. We sketch the proof following the note.
Let $\psi$ be a smooth function supported on [1,1][-1,1] such that $\sum_{k\in\mathbb{Z}}\psi(\cdot+k)=1$ on $\mathbb{R}$. Consider the partition of $\mathbb{R}$ by intervals $J$ of length $R^{-1}$ centered at $c_{J}\in R^{-1}\mathbb{Z}$. Let $\psi_{J}(\xi)=\psi(R(\xi-c_{J}))$ so that $\sum_{J}\psi_{J}=1$. We then decompose $f=\sum_{J}f_{J}$, where $\widehat{f}_{J}=\widehat{f}\psi_{J}$.
Fix $\tilde{\psi}\in C_{c}^{\infty}(\mathbb{R})$ such that $\tilde{\psi}\psi=\psi$. We may write $\mathbf{U}_{R^{2}}f_{J}(x,t)=K_{J}^{t}\ast f_{J}(x)$, where
$K_{J}^{t}(x)=(2\pi)^{-1}\eta(R^{-2}t)\int e^{ix\xi+it\xi^{2}}\tilde{\psi}(2R(\xi-c_{J}))\,d\xi.$
By changing variables $\xi\rightarrow R^{-1}\xi+c_{J}$, we observe that
$|K_{J}^{t}(x)|=(2\pi R)^{-1}\Big|\eta(R^{-2}t)\int e^{iR^{-1}(x+2tc_{J})\xi+iR^{-2}t\xi^{2}}\tilde{\psi}(2\xi)\,d\xi\Big|.$
Integration by parts yields the decay estimate
(45) $|K_{J}^{t}(x)|\leq C_{N}R^{-1}(1+R^{-1}|x+2tc_{J}|+R^{-2}|t|)^{-N},\quad N\geq 1,$
which, in particular, implies $\|K_{J}^{t}\|_{L_{x}^{1}}\lesssim 1$. Consequently, by the Cauchy-Schwarz inequality, we have $|\mathbf{U}_{R^{2}}f_{J}(x,t)|^{2}\lesssim|K_{J}^{t}|*|f_{J}|^{2}(x).$ For $q=(p/2)^{\prime}$, let $g\in L^{q}(\mathbb{R}^{2})$ with $\|g\|_{L^{q}}=1$. Thus we obtain
$\int\sum_{J}|\mathbf{U}_{R^{2}}f_{J}(x,t)|^{2}g(x,t)\,dxdt\lesssim\int\sum_{J}|f_{J}(y)|^{2}\mathfrak{M}g(y)\,dy,$
where $\mathfrak{M}g$ is defined by
$\mathfrak{M}g(y)=\sup_{J}\int|K_{J}^{t}(x-y)g(x,t)|\,dxdt,\quad y\in\mathbb{R}.$
By duality and Hölder, we find that
$\Big\|\big(\sum_{J}|\mathbf{U}_{R^{2}}f_{J}|^{2}\big)^{1/2}\Big\|_{L^{p}}^{2}\lesssim\big\|\big(\sum_{J}|f_{J}|^{2}\big)^{1/2}\big\|_{p}^{2}\,\,\big\|\mathfrak{M}\big\|_{L^{q}\to L^{q}}.$
By the one-dimensional analogue of (12), for $2\leq p\leq\infty$, $\big\|\big(\sum_{J}|f_{J}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R})}\lesssim\|f\|_{L^{p}(\mathbb{R})}.$ Thus in order to establish (44), it remains to prove that for $2\leq p\leq 4$,
(46) $\big\|\mathfrak{M}\big\|_{L^{q}\to L^{q}}\lessapprox R^{2\cdot\frac{2}{p}},\;\;q=\Big(\frac{p}{2}\Big)^{\prime}.$
To show this, for $w\in[-1,1]$ we set $T_{w}=\{(x,t)\in\mathbb{R}^{2}:|x+2tw|\leq R^{-1},~|t|\leq 1\},$ and define the Nikodym maximal function by
$\mathfrak{N}g(y):=\sup_{w\in[-1,1]}\frac{1}{|T_{w}|}\int_{T_{w}+(y,0)}|g(x,t)|\,dxdt,\quad y\in\mathbb{R}.$
The Nikodym maximal function satisfies the following bounds; see e.g. [33, 41].
Proposition 5.3.
For $2\leq q\leq\infty$ and $R\geq 1$, we have
$\|\mathfrak{N}g\|_{L^{q}(\mathbb{R})}\lessapprox\|g\|_{L^{q}(\mathbb{R})}.$
In view of (45), the operator $\mathfrak{M}$ can be dominated by a rescaled version of $\mathfrak{N}$. Indeed, after the rescaling $y\rightarrow R^{2}y$ and $(x,t)\rightarrow(R^{2}x,R^{2}t)$ and noting $|T_{w}|=R^{-1}$, we have
$\|\mathfrak{M}\|_{q\rightarrow q}\lesssim R^{2}R^{-\frac{2}{q}}\|\mathfrak{N}\|_{q\rightarrow q}\lessapprox R^{\frac{2}{q^{\prime}}}=R^{\frac{4}{p}},$
which verifies (46).
6. Examples for lower bounds
In this section, we discuss the lower bounds for the regularity $\zeta$ and $\gamma$ in 2.3, Theorem 2.4, and Theorem 2.5.
6.1. Lower bounds for $\zeta$ in 2.3
We show that (17) holds only if
(47) $\zeta\geq\zeta(\alpha,p):=\max\Big(\frac{1}{2}-\frac{1}{p},~\min\Big(\frac{\alpha}{2p},\frac{2\alpha-1}{2p}\Big)\Big).$
$(i)$ Proof of $\zeta\geq\frac{1}{2}-\frac{1}{p}$. To show this, we use an example from [35]. Let $\psi\in C_{c}^{\infty}([1/4,4])$ be such that $\psi=1$ on $[1/2,2]$. We take $\widehat{f}(\xi)=e^{-iR\xi^{2}}\psi(\xi)$ so that
$f(x)=(2\pi)^{-1}\int e^{ix\xi-iR\xi^{2}}\psi(\xi)\,d\xi.$
By integration by parts, we obtain $|f(x)|\lesssim(R+|x|)^{-N}$ for $|x|\geq 10R$ while the stationary phase method gives $|f(x)|\lesssim R^{-1/2}$ for $|x|\leq 10R$. Hence $\|f\|_{L^{p}}\lesssim R^{-\frac{1}{2}+\frac{1}{p}}.$
Note that
$|\mathbf{U}_{R}f(x)|\geq(2\pi)^{-1}\Big|\int e^{ix\xi+i(t-R)\xi^{2}}\psi(\xi)\,d\xi\Big|.$
In particular, one has $|\mathbf{U}_{R}f(x)|\gtrsim 1$ on the set
$F=\{(x,t)\in\mathbb{R}^{2}:|x|\leq c,~|t-R|\leq c\}$
for some small constant $c>0$. Thus
$\mu(F)^{\frac{1}{p}}\lesssim\big\|\mathbf{U}_{R}f\big\|_{L^{p}(\mathbb{R}\times[0,R],\mu)}\lesssim\langle\mu\rangle_{\alpha}^{\frac{1}{p}}R^{\zeta}R^{-\frac{1}{2}+\frac{1}{p}}$
follows. If we take $\mu=1_{F}(x,t)dxdt$, then $\langle\mu\rangle_{\alpha}\leq 1$ for any $\alpha\in[0,2]$. Since $\mu(F)\sim 1$, we obtain $\zeta\geq\frac{1}{2}-\frac{1}{p}$ as desired.
$(ii)$ Proof of $\zeta\geq\min\big(\frac{\alpha}{2p},\frac{2\alpha-1}{2p}\big)$. For a smooth function $\psi$ as above, choose $g$ such that $\widehat{g}(\xi)=\psi(R^{\frac{1}{2}}(\xi+1))$. Then $\|g\|_{L^{p}}\lesssim R^{-\frac{1}{2}+\frac{1}{2p}}$.
By the change of variable $\xi\rightarrow R^{-\frac{1}{2}}\xi-1$, ignoring the extra oscillatory factor independent of $\xi$, we obtain
$\big|\mathbf{U}_{R}g(x)\big|=(2\pi)^{-1}\Big|\int e^{ix\xi+it\xi^{2}}\psi(R^{\frac{1}{2}}(\xi+1))\,d\xi\Big|=(2\pi)^{-1}R^{-\frac{1}{2}}\Big|\int e^{iR^{-1/2}(x-2t)\xi+itR^{-1}\xi^{2}}\psi(\xi)\,d\xi\Big|.$
Thus $|\mathbf{U}_{R}g(x)|\gtrsim R^{-\frac{1}{2}}$ for $(x,t)\in G$ where
(48) $G=\{(x,t):|x-2t|\leq cR^{1/2},~|t|\leq R\}$
for a small constant $c>0$. Thus
(49) $R^{-\frac{1}{2}}\mu(G)^{\frac{1}{p}}\lesssim\big\|e^{it\partial_{x}^{2}}g\big\|_{L^{p}(\mathbb{R}\times[0,R],\mu)}\lesssim\langle\mu\rangle_{\alpha}^{1/p}R^{\zeta-\frac{1}{2}+\frac{1}{2p}}.$
For a set $G$ given in (48), cover $G$ by union of $O(R^{\frac{1}{2}})$ disjoint balls of radius $R^{\frac{1}{2}}$. When $\alpha\in[1,2]$ we obtain
$|G\cap B_{\rho}(z)|\lesssim\begin{cases}\rho^{2}\leq R^{\frac{2-\alpha}{2}}\rho^{\alpha},\quad~&\rho\in[1,R^{1/2}],\\[3.44444pt] R^{\frac{1}{2}}\rho\leq R^{\frac{2-\alpha}{2}}\rho^{\alpha},\quad~&\rho\in[R^{1/2},\infty).\end{cases}$
When $\alpha\in[0,1]$, we have
$|G\cap B_{\rho}(z)|\lesssim\begin{cases}\rho^{2}\leq R^{\frac{2-\alpha}{2}}\rho^{\alpha}\leq R^{\frac{3}{2}-\alpha}\rho^{\alpha},\quad~&\rho\in[1,R],\\[3.44444pt] R^{\frac{1}{2}}\rho\leq R^{\frac{3}{2}-\alpha}\rho^{\alpha},\quad~&\rho\in[R^{1/2},R],\\[3.44444pt] R^{\frac{3}{2}}\leq R^{\frac{3}{2}-\alpha}\rho^{\alpha},\quad~&\rho\in[R,\infty).\end{cases}$
We now take $\mu=\min\big(R^{\frac{\alpha-2}{2}},R^{\alpha-\frac{3}{2}}\big)1_{G}(x,t)dxdt$. Then $\langle\mu\rangle_{\alpha}\leq 1$ follows. Since $|G|\sim R^{\frac{3}{2}}$, we have $\mu(G)\sim\min\big(R^{\frac{\alpha+1}{2}},R^{\alpha}\big)$. Thus (49) yields the lower bounds $\zeta\geq\min\big(\frac{\alpha}{2p},\frac{2\alpha-1}{2p}\big)$.
6.2. Lower bounds for Theorem 2.4
To get lower bounds for $\gamma$ in Theorem 2.4, we rescale the estimates in Section 6.1 by following the argument in Section 5.1.
Suppose
(50) $\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1],\,d\mu)}\;\leq\;C\,[\mu]_{\beta,\text{par}}^{1/p}\,R^{\gamma}\|f\|_{L^{p}(\mathbb{R})}$
holds for all positive measure $\mu$ satisfying $\mu]_{\beta,\text{par}}\leq 1$ where $\widehat{f}$ is Fourier supported on $[R/2,R]$. For a given $f$, we set $f_{R}=f(R^{-1}\cdot)$ so that $\widehat{f_{R}}$ is supported on $[1/2,1]$ and $\|f_{R}\|_{p}=R^{\frac{1}{p}}\|f\|_{p}$. Similarly as before in Section 5.1, we define a positive measure $\mu_{R}$ by $\mu_{R}(E)=\int 1_{E}(Rx,R^{2}t)d\mu(x,t)$ for any $E\subset\mathbb{R}^{2}$. Then by (42), we have
$\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1],\mu)}=\big\|\mathbf{U}_{R^{2}}f_{R}\big\|_{L^{p}(\mathbb{R}\times[0,R^{2}],\mu_{R})}.$
Next, for a given positive Borel measure $\mu$ satisfying $\mu]_{\beta,\text{par}}\leq 1$, Lemma 5.1 provides
$\langle R^{\beta}\mu_{R}\rangle_{\theta}\lesssim 1$
where $\theta=\frac{\beta+1}{2}$ when $\beta\in[1,3]$, and $\theta=\beta$ when $\beta\in[0,1]$. By applying (50), we have
(51) $\big\|\mathbf{U}_{R^{2}}f_{R}\big\|_{L^{p}(\mathbb{R}\times[0,R^{2}],R^{\beta}\mu_{R})}\lesssim R^{\frac{\beta}{p}+\gamma}\|f\|_{L^{p}}=R^{\frac{\beta}{p}+\gamma-\frac{1}{p}}\|f_{R}\|_{L^{p}}.$
By the discussion in Section 6.1, we prove that if (51) holds, then $\frac{\beta}{p}+\gamma-\frac{1}{p}\geq 2\zeta(\theta,p)$ which is defined by (47). Equivalently, (51) holds only if
$\gamma\geq\begin{cases}2\max\big(\frac{1}{2}-\frac{1}{p},~\frac{\theta}{2p}\big)+\frac{1-\beta}{p},\quad~&\theta\in[1,2],\\[4.30554pt] 2\max\big(\frac{1}{2}-\frac{1}{p},~\frac{2\theta-1}{2p}\big)+\frac{1-\beta}{p},\quad~&\theta\in[0,1]\end{cases}$
where $\theta=\frac{\beta+1}{2}$ when $\beta\in[1,3]$, and $\theta=\beta$ when $\beta\in[0,1]$. Hence,
$\gamma\geq\begin{cases}\max\!\big\{1-\frac{\beta+1}{p},\,\frac{3-\beta}{2p}\big\},&\beta\in[1,3],\\[4.30554pt] \max\!\big\{1-\frac{\beta+1}{p},\,\frac{\beta}{p}\},&\beta\in[0,1].\end{cases}$
6.3. Proof of Theorem 2.5
We now discuss the lower bounds for $\gamma$ of Theorem 2.5. Suppose
(52) $\big\|e^{it\partial_{x}^{2}}f\big\|_{L^{p}(\mathbb{R}\times[0,1],\,d\mu)}\;\leq\;C\,[\mu]_{\alpha}^{1/p}\,R^{\gamma}\|f\|_{L^{p}(\mathbb{R})}$
holds. Let $\mu$ be a positive measure satisfying $\mu]_{\alpha}\leq 1$ for $\alpha\in[0,2]$. By Lemma 5.1, we have
$\langle R^{\theta}\mu_{R}\rangle_{\alpha}\lesssim 1$
for $\theta=2\alpha-1$ (when $\alpha\in[1,2]$) and $\theta=\alpha$ (when $\alpha\in[0,1]$). As before, if (52) holds, then
(53) $\big\|\mathbf{U}_{R^{2}}f_{R}\big\|_{L^{p}(\mathbb{R}\times[0,R^{2}],R^{\theta}\mu_{R})}\lesssim R^{\frac{\theta}{p}+\gamma}\|f\|_{L^{p}}=R^{\frac{\theta}{p}+\gamma-\frac{1}{p}}\|f_{R}\|_{L^{p}}.$
By the discussion in Section 6.1, (53) holds only if $\frac{\theta}{p}+\gamma-\frac{1}{p}\geq 2\zeta(\alpha,p)$. Equivalently, we have
$\gamma\geq\begin{cases}2\max\big(\frac{1}{2}-\frac{1}{p},~\frac{\alpha}{2p}\big)+\frac{1-(2\alpha-1)}{p},\quad~&\theta\in[1,2],\\[4.30554pt] 2\max\big(\frac{1}{2}-\frac{1}{p},~\frac{2\alpha-1}{2p}\big)+\frac{1-\alpha}{p},\quad~&\theta\in[0,1].\end{cases}$
This yields
$\gamma\geq\begin{cases}\max\!\big(1-\frac{2\alpha}{p},\,\frac{2-\alpha}{p}\big),&\alpha\in[1,2],\\[3.44444pt] \max\!\big(1-\frac{\alpha+1}{p},\,\frac{\alpha}{p}\big),&\alpha\in[0,1].\end{cases}$
Appendix A Proof of Lemma 4.1
In this section, we prove Lemma 4.1, motivated by the argument in [5]. The following elementary lemma will be used repeatedly.
Lemma A.1.
Let $\{a_{i}\}_{i\in I}$ be a sequence of non-negative real number indexed by a finite set $I$. For each $i\in I$, let $I_{i}\subset I$ be a subset containing $i$ such that $|I_{i}|\leq C_{1}$ for all $i\in I$ for some constant $C_{1}\in\mathbb{N}$. Then there exists $C=C(C_{1},p)$ such that for $p\geq 1$,
$\Big(\sum_{i\in I}a_{i}\Big)^{p}\leq C\Big(\max_{i\in I}a_{i}^{p}+(\#I)^{p}\max_{\begin{subarray}{c}i\in I,\\ j\notin I_{i}\end{subarray}}a_{i}^{\frac{p}{2}}a_{j}^{\frac{p}{2}}\Big).$
Proof.
Let $\in I$ be the indices for which $a_{*}=\max_{i\in I}a_{i}$ and $a_{**}=\max_{j\notin I_{*}}a_{j}$. Then, for any $i\notin I_{*}$ we have $a_{i}\leq a_{*}^{1/2}a_{**}^{1/2}$. Therefore,
$\sum_{i\in I}a_{i}$ $=\sum_{i\in I_{*}}a_{i}+\sum_{i\notin I_{*}}a_{i}\leq C_{1}a_{*}+(\#I)a_{*}^{1/2}a_{**}^{1/2}.$
Taking $p$-th power and using the inequality $(x+y)^{p}\lesssim_{p}x^{p}+y^{p}$ for $x,y>0$, we obtain the desired bound. ∎
Now we prove Lemma 4.1.
Proof of Lemma 4.1.
Let $R\geq 1$. Recall that $K$ is a dyadic number chosen so that
$1\leq K\leq\dots\leq K^{m}\sim R^{1/2}.$
Let $\mathcal{T}_{0}=\{\tau_{0}\}$ denote collection of the unit cube covering the parabola $\mathcal{P}$.
At the first stage, we decompose $\tau_{0}$ into a collection $\mathcal{T}_{1}=\mathcal{T}_{1}(\tau_{0})$ of $K^{-1}\times K^{-2}$ boxes $\tau_{1}$ covering the $K^{-2}$-nbd of the parabola. For each $\tau_{1}\in\mathcal{T}_{1}$, let $\mathcal{T}_{2}(\tau_{1})$ be a collection of $K^{-2}\times K^{-4}$ boxes $\tau_{2}\subset\tau_{1}$ covering $K^{-4}$-nbd of $\mathcal{P}$, and define $\mathcal{T}_{2}=\cup_{\tau_{1}\in\mathcal{T}_{1}}\mathcal{T}_{2}(\tau_{1})$.
Proceeding inductively, for $2\leq j\leq m$, we define $\mathcal{T}_{j}(\tau_{j-1})$ as the collection of boxes $\tau_{j}\subset\tau_{j-1}$ of dimension $K^{-j}\times K^{-2j}$ covering $K^{-2j}$-neighborhood of $\mathcal{P}$ and $\mathcal{T}_{j}$ similarly. Finally, denote $\mathcal{T}_{m}=\{\theta\}$ and for each $j=1,\dots,m$ set
$f_{\tau_{j}}=\sum_{\theta\subset\tau_{j}}f_{\theta}.$
For each $\tau_{1}\in\mathcal{T}_{1}$, let
$\mathcal{N}_{1}(\tau_{1})=\{\tau_{1}^{\prime}\in\mathcal{T}_{1}:~\tau_{1}^{\prime}\cap 2\tau_{1}\neq\emptyset\}.$
It is clear that $\mathcal{N}_{1}(\tau_{1})$ consists of only $O(1)$ many elements, and if $\tau_{1}^{\prime}\notin\mathcal{N}_{1}(\tau_{1})$ then $d(\tau_{1},\tau_{1}^{\prime})\geq 1/K$. Since $\mathcal{T}_{1}$ is covered by such neighborhood $\mathcal{N}_{1}(\tau_{1})$, applying Lemma A.1 to $|f|^{p}\leq(\sum_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}|)^{p}$, and using that $\mathcal{T}_{1}\lesssim K$, we have
$|f(x)|^{p}\leq C\max_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}(x)|^{p}+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq\frac{1}{2K}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}$
for some absolute constant $C$.
Applying Lemma A.1 again to the first term $|f_{\tau_{1}}|^{p}=|\sum_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}f_{\tau_{2}}|^{p}$, we get
$|f(x)|^{p}$ $\leq C^{2}\max_{\tau_{1}}\max_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}|f_{\tau_{2}}(x)|^{p}+C^{2}K^{p}\max_{\tau_{1}}\max_{\begin{subarray}{c}\tau_{2},\tau_{2}^{\prime}\in\mathcal{T}_{2}(\tau_{1});\\ d(\tau_{2},\tau_{2}^{\prime})\geq\frac{1}{2K^{2}}\end{subarray}}|f_{\tau_{2}}(x)|^{\frac{p}{2}}|f_{\tau_{2}^{\prime}}(x)|^{\frac{p}{2}}$
$\qquad\quad+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq\frac{1}{2K}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}.$
Continuing in this manner, with $\mathcal{T}:=\mathcal{T}_{m}$, we get
$\begin{split}|f(x)|^{p}&\leq C^{m}\max_{\theta\in\mathcal{T}}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\max_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq\frac{1}{2K^{j}}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}\\ &\leq C^{m}\sum_{\theta\in\mathcal{T}}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\sum_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq\frac{1}{2K^{j}}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}.\end{split}$
This completes the proof of Lemma 4.1. ∎
Acknowledgement
J. Kim was supported in part by grants from the Research Grants Council of the Hong Kong Administrative Region, China (Project No. CityU 21309222 and CityU 11308924). H. Ko was supported in part by the NRF (Republic of Korea) grant RS-2024-00339824, G-LAMP RS-2024-00443714, JBNU research funds for newly appointed professors in 2024, and NRF2022R1I1A1A01055527. We thank the organizers of the Special Topic School “Maximal Operators and Applications” and the Hausdorff Center for Mathematics, where our collaboration began. We also thank Sanghyuk Lee for helpful discussions. The second author would like to thank Seheon Ham for previous collaboration related to this topic.

### v2

@@
 Abstract.
-In this paper, we extend Fefferman’s classical square function estimate for the parabola to a weighted setting.
+In this paper, we extend the Córdoba–Fefferman square function estimate for the parabola to a weighted setting.
 Our weighted square function estimate is derived from a weighted wave envelope estimate for the parabola.
@@
 (1) $\|f\|_{L^{4}(\mathbb{R}^{2})}\leq C\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{4}(\mathbb{R}^{2})}.$
-This inequality relies on the geometric observation by Fefferman [14] that the algebraic differences $\theta-\theta^{\prime}$ overlap only finitely often as $\theta\neq\theta^{\prime}$ vary.
+This inequality relies on the geometric observation by Fefferman [16] that the algebraic sums $\theta+\theta^{\prime}$ overlap only finitely often as $\theta$ and $\theta^{\prime}$ vary.
 See also [19, 32, 22] for extensions to non-degenerate curves in higher dimensions.
@@
 The sharp square function estimate (1) is known to imply sharp results for the Kakeya maximal function, the Bochner-Riesz multipliers, the Fourier restriction operator, and local smoothing estimates for the Schrödinger equation; see [6, 43] and references therein.
-In higher dimensions, it is conjectured that (1) holds with $L^{4}(\mathbb{R}^{2})$ replaced by $L^{\frac{2d}{d-1}}(\mathbb{R}^{d})$, which remains wide open.
+For the paraboloid in higher dimensions, it is conjectured that (1) holds with $L^{4}(\mathbb{R}^{2})$ replaced by $L^{\frac{2d}{d-1}}(\mathbb{R}^{d})$, which remains wide open.
 Let $H:\mathbb{R}^{2}\to[0,\infty)$ be a bounded function on $\mathbb{R}^{2}$. The main goal of this paper is to establish weighted square function estimates of the form
@@
 (4) $\|f\|_{L^{p}(Hdx)}\lessapprox\big(\max_{U\in\mathbb{U}}\kappa_{p,H}(U)+R^{-100}\big)\big\|\big(\sum_{\theta}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{p}(\mathbb{R}^{2})}.$
-Here we mean by $A\lessapprox B$ an inequality of the form $A\leq C_{p}(\log R)^{O(1)}B$.
+Here we mean by $A\lessapprox B$ an inequality of the form $A\leq C_{\epsilon}R^{\epsilon}B$ for any $\epsilon>0$.
 The term $R^{-100}$ on the right-hand side of (4) is negligible for interesting weights $H$.
@@
 For $2\leq p<4$, the $H\equiv 1$ case of Theorem 1.1 recovers square function estimates due to Gan [17], where more general small cap square function estimates are established.
+The case $p=2$ of Theorem 1.1 is essentially contained in [8]; this was kindly pointed out to us by Tony Carbery (see also [7, pages 17–19]).
 Our weighted square function estimates are inspired by weighted decoupling inequalities for the paraboloids, which have been extensively studied in recent years and applied to problems such as the Falconer distance set conjecture and Bochner-Riesz means; see, e.g., [20, 13, 18, 28] and references therein.
@@
 Their wave envelope estimate for $\Gamma$ implies, among other consequences, the sharp $L^{4}$ square function estimate for $\Gamma$ and the sharp local smoothing estimate for the wave equation in 2+1 dimension.
-The $p=4$ and $H\equiv 1$ case of Theorem 1.4 recovers the $L^{4}$ wave envelope estimate [19, Equation (8)], which is implicit in [21].
+The $p=4$ and $H\equiv 1$ case of Theorem 1.4 recovers the $L^{4}$ wave envelope estimate [21, Equation (8)], which is implicit in [23].
 We note that [21] established more refined versions of the $L^{4}$ wave envelope estimates, termed amplitude-dependent wave envelope estimates, for both the parabola $\mathcal{P}$ and the cone $\Gamma$.
@@
 We summarize here the notations that will be used frequently throughout the paper.
-$\bullet$ We write $A\lessapprox B$ to denote an inequality of the form $A\leq C_{p}(\log R)^{O(1)}B$ for $R>1$ where $C_{p}$ is a constant depending only on $p$.
+$\bullet$ We write $A\lessapprox B$ to denote an inequality of the form $A\leq C_{\epsilon}R^{\epsilon}B$ for any $\epsilon>0$ where $C_{\epsilon}$ is an absolute constant.
 $\bullet$ We denote by $B_{\rho}(z)$ the ball of radius $\rho$ centered at $z\in\mathbb{R}^{2}$ and we simply write $B_{\rho}$ when centered at the origin.
@@
 (14) $\Big\|\big(\sum_{I}|\mathbf{U}_{R}f_{I}|^{2}\big)^{1/2}\Big\|_{L^{p}(\mathbb{R}^{2})}\lessapprox R^{\frac{1}{p}}\|f\|_{L^{p}(\mathbb{R}^{2})},\;\;2\leq p\leq 4.$
-A detailed proof of (14) can be found in a note by Yung [41, Proof of Theorem 2].
+A detailed proof of (14) can be found in a note by Yung [43, Proof of Theorem 2].
 It follows from a standard duality argument similar to the one used in [11] and [35].
@@
 Thus, $\langle H\rangle_{\alpha}\lesssim 1$. A similar computation shows that $\|H\|_{\infty}\lesssim 1$. Consequently, Theorem 2.2 and (5) give the desired estimate. ∎
-[…]
+2.3 can be obtained by interpolation between known $L^{2}$ and $L^{4}$ estimates.
+Indeed, when $p=2$, 2.3 recovers a bound due to Du and Zhang [14] (see also [42, 15]).
+The $p=4$ case of 2.3 can be deduced from a local smoothing estimate for the Schrödinger equation:
 (16) $\|e^{it\partial_{x}^{2}}f\|_{L^{p}(\mathbb{R}\times[0,1])}\leq C_{p,\gamma}\|f\|_{L^{p}_{\gamma}(\mathbb{R})},\;\;p\in(2,\infty)\;\text{and}\;\gamma>\max\Big(0,1-\frac{4}{p}\Big),$
@@
 $\nu((t-\rho,t+\rho))\leq C\rho^{\alpha},\quad(t,\rho)\in\mathbb{R}\times\mathbb{R}_{+}.$
-In [28, Theorem 1.6], it is shown that if $\gamma\geq-\frac{\alpha}{2}$, then $H^{\gamma}$$L_{t}^{2}(d\nu;L_{x}^{2}(-1,1))$ estimates holds for $e^{it\partial_{x}^{2}}f$.
+In [30, Theorem 1.6], it is shown that if $\gamma\geq-\frac{\alpha}{2}$, then $H^{\gamma}$$L_{t}^{2}(d\nu;L_{x}^{2}(-1,1))$ estimates holds for $e^{it\partial_{x}^{2}}f$.
 Moreover, optimal results in higher dimensions and weighted Strichartz-type estimates of the form $H^{\gamma}$$L_{t}^{q}(d\nu;L_{x}^{r})$ for fractional Schrödinger operators are established in [30] (see also [4] for analogous results for the wave operator).
@@
 4.2. Decomposition
-For $R\geq 1$, let $K$ be a dyadic number such that $K\sim\log R$.
-We choose $m\in\mathbb{N}$ such that $K^{m}\sim R^{1/2}$.
-We consider larger canonical covering $\tau$ such that $|\tau|=s$ for each scale
+Let $K\sim\log R$ be a dyadic number. Let $m\in\mathbb{N}$ be the integer such that $K^{m}\leq R^{1/2}<K^{m+1}$. We consider a canonical covering $\{\tau\}$ such that $|\tau|=s$ for each scale
 $s=K^{-j},\quad j=0,1,\cdots,m.$
-At the smallest scale for $j=m$, we obtain the collection of blocks $\{\theta\}$ of size $|\theta|=K^{-m}\sim R^{-1/2}.$ For each block $\tau$ at an intermediate scale, we define
+For each block $\tau$ at an intermediate scale, we define
 $f_{\tau}=\sum_{\theta\subset\tau}f_{\theta}.$
-We use the following pointwise estimate from [10, Section 5].
+We use the following pointwise estimate from [12, Section 5].
 Lemma 4.1.
-For any $x\in\mathbb{R}^{2}$, there is an absolute constant $C>0$ (independent of $x$) such that
-(32) $|f(x)|^{p}\leq C^{m}\sum_{\theta}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{R^{-1/2}<s\leq 1}\sum_{|\tau|=s}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-1}s\\ d(\tau_{1},\tau_{2})\geq\frac{1}{2}K^{-1}s\end{subarray}}|f_{\tau_{1}}(x)f_{\tau_{2}}(x)|^{\frac{p}{2}},$
+For any $x\in\mathbb{R}^{2}$, we have the pointwise bound
+(32) $|f(x)|^{p}\lessapprox\sum_{|\theta|=R^{-1/2}}|f_{\theta}(x)|^{p}+\sum_{1\leq j\leq m}\sum_{|\tau|=K^{-j+1}}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-j}\\ d(\tau_{1},\tau_{2})\geq K^{-j}\end{subarray}}|f_{\tau_{1}}(x)f_{\tau_{2}}(x)|^{\frac{p}{2}},$
 where $d(\tau_{1},\tau_{2})$ denotes the distance between $\tau_{1}$ and $\tau_{2}$.
@@
 By integrating (32) over $Y$, we get
-(33) $\|f\|_{L^{p}(Y)}^{p}\!\leq\!C^{m}\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}+C^{m}K^{p}\sum_{R^{-1/2}<s\leq 1}\sum_{|\tau|=s}\sum_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-1}s\\ d(\tau_{1},\tau_{2})\geq\frac{1}{2}K^{-1}s\end{subarray}}\!\!\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p}.$
+(33) $\|f\|_{L^{p}(Y)}^{p}\lessapprox\sum_{\theta}\|f_{\theta}\|_{L^{p}(Y)}^{p}+\sum_{1\leq j\leq m}\sum_{|\tau|=K^{-j+1}}\max_{\begin{subarray}{c}\tau_{1},\tau_{2}\subset\tau\\ |\tau_{1}|=|\tau_{2}|=K^{-j}\\ d(\tau_{1},\tau_{2})\geq K^{-j}\end{subarray}}\!\!\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p},$
+where we have used the fact that, for each fixed $\tau$, there are at most $K^{O(1)}\lessapprox 1$ pairs $\tau_{1},\tau_{2}\subset\tau$.
 By the computation (9), the first term on the right-hand side of (33) is bounded by
@@
 $\big\||f_{\tau_{1}}|^{\frac{1}{2}}|f_{\tau_{2}}|^{\frac{1}{2}}\big\|_{L^{p}(Y)}^{p}\lessapprox\sum_{U\in\mathbb{U}_{\tau}}\kappa_{p,Y}(U)^{p}|U|^{1-\frac{p}{2}}\big\|\big(\sum_{\theta\subset\tau}|f_{\theta}|^{2}\big)^{1/2}\big\|_{L^{2}(w_{U})}^{p}.$
-Since for each fixed $\tau$ there are only $K^{O(1)}$ (and hence $\lesssim\log R$) such pairs $\tau_{1},\tau_{2}\subset\tau$, this logarithmic factor can be absorbed into the $O((\log R)^{O(1)})$ loss.
-Consequently, the bilinear term in the sum in (33) is bounded by the right-hand side of (8).
+This estimate bounds the bilinear terms on the right-hand side of (33) by the right-hand side of (8).
 4.3. Proof of 4.2
@@
 where $w_{B}$ is a $L^{\infty}$-normalized weight which decays rapidly away from $B$.
-The proof of Theorem 4.4 is standard; see, for example, [38] or [29, Lemma 2.4].
+The proof of Theorem 4.4 is standard; see, for example, [40] or [31, Lemma 2.4].
 We now turn to the proof of Lemma 4.3.
@@
 Proof of Lemma 4.1.
-Let $R\geq 1$. Recall that $K$ is a dyadic number chosen so that
-$1\leq K\leq\dots\leq K^{m}\sim R^{1/2}.$
+Recall that $K\sim\log R$ is a dyadic number and $m$ is the integer such that
+$1\leq K\leq\dots\leq K^{m}\leq R^{1/2}<K^{m+1}.$
 Let $\mathcal{T}_{0}=\{\tau_{0}\}$ denote collection of the unit cube covering the parabola $\mathcal{P}$.
@@
 Since $\mathcal{T}_{1}$ is covered by such neighborhood $\mathcal{N}_{1}(\tau_{1})$, applying Lemma A.1 to $|f|^{p}\leq(\sum_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}|)^{p}$, and using that $\mathcal{T}_{1}\lesssim K$, we have
-$|f(x)|^{p}\leq C\max_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}(x)|^{p}+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq\frac{1}{2K}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}$
+$|f(x)|^{p}\leq C\max_{\tau_{1}\in\mathcal{T}_{1}}|f_{\tau_{1}}(x)|^{p}+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq K^{-1}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}$
 for some absolute constant $C$.
 Applying Lemma A.1 again to the first term $|f_{\tau_{1}}|^{p}=|\sum_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}f_{\tau_{2}}|^{p}$, we get
-$|f(x)|^{p}$ $\leq C^{2}\max_{\tau_{1}}\max_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}|f_{\tau_{2}}(x)|^{p}+C^{2}K^{p}\max_{\tau_{1}}\max_{\begin{subarray}{c}\tau_{2},\tau_{2}^{\prime}\in\mathcal{T}_{2}(\tau_{1});\\ d(\tau_{2},\tau_{2}^{\prime})\geq\frac{1}{2K^{2}}\end{subarray}}|f_{\tau_{2}}(x)|^{\frac{p}{2}}|f_{\tau_{2}^{\prime}}(x)|^{\frac{p}{2}}$
-$\qquad\quad+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq\frac{1}{2K}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}.$
-Continuing in this manner, with $\mathcal{T}:=\mathcal{T}_{m}$, we get
-$\begin{split}|f(x)|^{p}&\leq C^{m}\max_{\theta\in\mathcal{T}}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\max_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq\frac{1}{2K^{j}}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}\\ &\leq C^{m}\sum_{\theta\in\mathcal{T}}|f_{\theta}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\sum_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq\frac{1}{2K^{j}}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}.\end{split}$
-This completes the proof of Lemma 4.1. ∎
+$|f(x)|^{p}$ $\leq C^{2}\max_{\tau_{1}}\max_{\tau_{2}\in\mathcal{T}_{2}(\tau_{1})}|f_{\tau_{2}}(x)|^{p}+C^{2}K^{p}\max_{\tau_{1}}\max_{\begin{subarray}{c}\tau_{2},\tau_{2}^{\prime}\in\mathcal{T}_{2}(\tau_{1});\\ d(\tau_{2},\tau_{2}^{\prime})\geq K^{-2}\end{subarray}}|f_{\tau_{2}}(x)|^{\frac{p}{2}}|f_{\tau_{2}^{\prime}}(x)|^{\frac{p}{2}}$
+$\qquad\quad+CK^{p}\max_{\begin{subarray}{c}\tau_{1},\tau_{1}^{\prime}\in\mathcal{T}_{1};\\ d(\tau_{1},\tau_{1}^{\prime})\geq K^{-1}\end{subarray}}|f_{\tau_{1}}(x)|^{\frac{p}{2}}|f_{\tau_{1}^{\prime}}(x)|^{\frac{p}{2}}.$
+Continuing in this manner, we get
+$\begin{split}|f(x)|^{p}&\leq C^{m}\max_{\tau_{m}\in\mathcal{T}_{m}}|f_{\tau_{m}}(x)|^{p}+C^{m}K^{p}\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\max_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq K^{-j}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}}\\ &\lessapprox\sum_{|\theta|=R^{-1/2}}|f_{\theta}(x)|^{p}+\sum_{j=1}^{m}\sum_{\tau_{j-1}\in\mathcal{T}_{j-1}}\sum_{\begin{subarray}{c}\tau_{j},\tau_{j}^{\prime}\in\mathcal{T}_{j}(\tau_{j-1});\\ d(\tau_{j},\tau_{j}^{\prime})\geq K^{-j}\end{subarray}}|f_{\tau_{j}}(x)|^{\frac{p}{2}}|f_{\tau_{j}^{\prime}}(x)|^{\frac{p}{2}},\end{split}$
+where we have used $C^{m}K^{O(1)}\lessapprox 1$ together with the fact that each $\tau_{m}\in\mathcal{T}_{m}$ contains $O(K)$ elements $\theta$ of size $|\theta|=R^{-1/2}$.
+This completes the proof of Lemma 4.1.
+∎
 Acknowledgement
@@
 The second author would like to thank Seheon Ham for previous collaboration related to this topic.
+We are grateful to Tony Carbery for bringing the references [8, 7] to our attention and for pointing out the overlap between their results and the case $p=2$ of Theorem 1.1.